### v1

# Formal Foundations for Controlled Stochastic Activity Networks

Abstract
We introduce Controlled Stochastic Activity Networks (Controlled SANs), a formal extension of classical Stochastic Activity Networks that integrates explicit control actions into a unified semantic framework for modeling distributed real-time systems. Controlled SANs systematically capture dynamic behavior involving nondeterminism, probabilistic branching, and stochastic timing, while enabling policy-driven decision making within a rigorous mathematical setting.
We develop a hierarchical, automata-theoretic semantics for Controlled SANs that encompasses nondeterministic, probabilistic, and stochastic models in a uniform manner. A structured taxonomy of control policies—ranging from memoryless and finite-memory strategies to computationally augmented policies—is formalized, and their expressive power is characterized through associated language classes. To support model abstraction and compositional reasoning, we introduce behavioral equivalences including bisimulation and stochastic isomorphism.
Controlled SANs generalize classical frameworks such as continuous-time Markov decision processes (CTMDPs), providing a rigorous foundation for the specification, verification, and synthesis of dependable systems operating under uncertainty. This framework enables both quantitative and qualitative analysis, advancing the design of safety-critical systems where control, timing, and stochasticity are tightly coupled.
keywords:
Controlled stochastic activity networks , controlled automata , probabilistic automata , continuous-time Markov decision processes , formal verification , reinforcement learning , safety-critical systems
journal: Information and Computation
1 Introduction
Safety-critical systems—such as those used in avionics, autonomous vehicles, medical devices, and industrial automation—are characterized by stringent correctness requirements, continuous operation, and the need to make timely decisions under uncertainty. These systems typically involve a distributed architecture with interacting components that operate concurrently and must respond predictably to a combination of real-time stimuli, probabilistic events, and external commands. Any error in their behavior can have catastrophic consequences, making formal modeling and verification indispensable to their development and deployment [hoare1969axiomatic, lamport1977proving, newcombe2015aws, woodcock2009survey, seshia2022cyberphysical].
To analyze and synthesize such systems, we require modeling frameworks that are expressive enough to capture their complex dynamic behavior, yet structured enough to support rigorous reasoning and tool-based analysis. Classical paradigms—such as finite-state machines [Sipser2013], timed automata [alur1994theory], Petri nets [Peterson1981], and stochastic process algebras [hillston1996compositional, hermanns2002interactive]—have each contributed key insights in this direction. In particular, Stochastic Activity Networks (SANs) [movaghar1984stochastic, meyer1985stochastic, movaghar2001stochastic] generalize timed Petri nets by associating activities with probabilistic durations and gate-driven transitions. SANs are equipped with intuitive graphical notations and support compositionality, making them effective for performance and dependability analysis.
Despite these strengths, SANs fall short in contexts that demand explicit control, i.e., decision logic that governs how the system evolves in response to observed events, internal states, or policies. In safety-critical domains, such control is pervasive, ranging from real-time schedulers and fault-tolerance strategies to adaptive control loops and fail-safe mechanisms. However, traditional SANs lack the semantic machinery to encode and reason about such decision-driven behavior. This limitation motivates the development of models that integrate control theory with stochastic modeling and formal semantics.
In this paper, we introduce Controlled Stochastic Activity Networks (Controlled SANs), a formal extension of SANs that incorporates control actions and policies into the modeling process. Our goal is to provide a unified formalism for modeling and analyzing systems where control decisions interact with stochastic and timed behaviors. The resulting models support analysis tasks essential to safety-critical systems, such as controller synthesis, reachability under uncertainty, risk evaluation, and reward optimization, within a compositional and mathematically rigorous framework.
Motivation from Safety-Critical Domains
Consider an autonomous drone navigating an urban environment. Its mission requires sensing obstacles, adjusting altitude and velocity in real-time, avoiding collisions, and ensuring safe return. Each of these behaviors involves uncertain timings (e.g., sensor latencies), probabilistic outcomes (e.g., wind disturbances), and control decisions (e.g., evasive maneuvers). Similarly, a medical infusion pump must adjust dosages based on monitored vitals, react to hardware faults, and enforce safety constraints on flow rates—all while managing concurrent activities with varied timing profiles. In both cases, a formal model must:
  • Represent concurrent activities with different timing semantics.
  • Express control actions that guide system transitions based on state, history, or external inputs.
  • Support probabilistic modeling of failures, delays, or random phenomena.
  • Enable quantitative reasoning about safety, liveness, and performance objectives.
Controlled SANs address these needs by extending SANs with policy-driven control, enabling formal reasoning about decision-making under uncertainty and time constraints. Through their integration of nondeterministic, probabilistic, and stochastic semantics, they provide a natural and expressive framework for modeling and analyzing safety-critical systems.
Overview of the Model and Contributions
We develop Controlled SANs through a layered approach that introduces increasing levels of semantic detail. At each level, we formalize system behavior using corresponding automata-theoretic models, define classes of control policies, and characterize the expressive power of the framework in terms of language recognition and bisimulation [Milner1989, baier2008principles].
1. Nondeterministic Semantics. At the foundational level, Controlled SANs extend classical activity networks by introducing control actions that mediate the outcome of timed activities. System evolution is governed by marking transitions induced by activity completions and gate functions. The behavior of such systems is captured by controlled automata, where transitions are labeled by activities and control actions. These automata define state-based semantics for Controlled SANs and allow for equivalence checking and reasoning about nondeterministic executions.
2. Probabilistic Semantics. To handle uncertainty in instantaneous transitions, we introduce controlled probabilistic activity networks, where the choice among enabled instantaneous activities in an unstable marking is made probabilistically. These are interpreted as controlled probabilistic automata, which generalize discrete-time Markov decision processes (DTMDPs) [puterman1994markov] with activity labels and control actions. This extension supports modeling fault injection, randomized algorithms, and probabilistic control behavior.
3. Stochastic Semantics. For real-time systems, we enrich the model with timing semantics by assigning probability distributions and enabling rates to timed activities. The resulting controlled stochastic activity networks are semantically equivalent to controlled stochastic automata, which model system dynamics as continuous-time stochastic processes governed by policies. In the special case where timing distributions are exponential, the models correspond to continuous-time Markov decision processes (CTMDPs) [puterman1994markov], enabling application of well-established analysis techniques.
Across all these models, we define a spectrum of policy types—including memoryless, history-dependent, finite-memory, stack-augmented, and tape-augmented—and classify their expressive power using formal language theory. This taxonomy offers a rigorous lens for assessing controller complexity and capability, particularly relevant to automated synthesis in safety-critical settings.
Theoretical and Practical Contributions
Our contributions in this paper include:
  1. 1.
    A unified formalism for modeling control-enhanced stochastic systems using Controlled SANs, encompassing nondeterministic, probabilistic, and stochastic dynamics.
  2. 2.
    Formal definitions and semantics for Controlled SANs based on automata-theoretic models, including controlled automata, controlled probabilistic automata, and controlled stochastic automata.
  3. 3.
    Introduction and classification of control policies by computational structure and memory, with a detailed study of their expressive power via language acceptance.
  4. 4.
    Bisimulation-based notions of equivalence, enabling compositional reasoning and model reduction while preserving behavioral semantics.
  5. 5.
    A formal correspondence between Controlled SANs and CTMDPs, demonstrating the compatibility of our model with established verification frameworks.
Through these results, Controlled SANs offer a rigorous foundation for the design, analysis, and synthesis of decision-making in stochastic real-time systems. They support both qualitative properties (e.g., safety, liveness, and controllability) and quantitative objectives (e.g., expected reward, risk bounds, or deadline satisfaction), making them suitable for a wide range of safety-critical domains.
While this paper focuses on the formal and semantic foundations of Controlled SANs rather than their algorithmic verification, the framework provides the theoretical groundwork for integrating temporal-logic model-checking techniques (such as CTL and LTL) in future research.
Organization of the Paper
The rest of the paper is organized as follows. Section 2 discusses the related work. Section 3 introduces the syntax and semantics of Controlled SANs in a nondeterministic setting, along with their interpretation via controlled automata. Section 4 extends the model to controlled probabilistic activity networks, discussing policy types and language classes. Section 5 presents the stochastic semantics, introduces controlled stochastic automata, and formalizes the transition to CTMDPs. Section 6 summarizes the key findings and outlines future directions, including verification techniques and safety assurance methods for real-world applications.
2 Related Work
The Controlled SAN framework introduced in this paper brings together several classical and modern strands of research in stochastic modeling, formal verification, and automata theory. In this section, we review the most relevant foundational and applied models that inform and contrast with our approach.
Petri Nets and SANs
Stochastic Activity Networks (SANs) [movaghar1984stochastic, meyer1985stochastic, movaghar2001stochastic] extend timed Petri nets by associating activities with probabilistic timing and gate-based control logic. SANs have been extensively applied to performance modeling of distributed and fault-tolerant systems. However, they lack mechanisms for encoding control policies—that is, dynamic selection among multiple transitions based on system state or history. Our Controlled SANs generalize SANs by integrating control actions and policy-based semantics, subsuming the original SAN model as a special case without decision-making.
The formal underpinnings of SANs are rooted in Petri net theory [Peterson1981], which has also been enriched through control-oriented extensions such as inhibitor arcs and priority transitions, and through the application of supervisory control theory to Petri nets [Ramadge1987, Cassandras2008]. These approaches often address control indirectly or externally. In contrast, Controlled SANs embed policy-driven behavior directly into the model’s semantics, enabling explicit and systematic reasoning about system evolution under control.
CTMDPs and GSMPs
Continuous-Time Markov Decision Processes (CTMDPs) [puterman1994markov] represent one of the canonical models for combining stochastic timing with nondeterministic control. They support optimal policy synthesis under quantitative objectives and are widely used in model checking and dependability analysis. Controlled SANs subsume CTMDPs in the special case where all timed activities follow exponential distributions and control actions determine probabilistic transitions. However, our framework allows richer structural and semantic features, including gate-driven transitions, compositionality, and generalized timing semantics.
Generalized Semi-Markov Processes (GSMPs) [Matthes1962, Schassberger1978] extend CTMDPs to support multiple concurrently enabled transitions with arbitrary distributions. While expressive, GSMPs lack an integrated control structure. Controlled SANs may be viewed as a control-augmented subclass of GSMPs, with a more explicit policy-theoretic semantics and better alignment with formal methods.
Timed Automata and Stochastic Hybrid Systems
Timed automata [alur1994theory] and their probabilistic and stochastic extensions [baier2008principles] offer a well-established formal framework for the specification and analysis of systems with real-time constraints. These models have been central to the development of verification tools such as UPPAAL and PRISM. However, traditional timed automata are clock-driven and best suited for deterministic or nondeterministic behaviors. Controlled SANs, in contrast, are activity-driven, using distributed timers with stochastic durations and gate-based enabling, which better captures concurrent probabilistic dynamics.
Stochastic Hybrid Systems (SHS)  [Cassandras2008, Tomlin1998] and Piecewise Deterministic Markov Processes (PDMPs)  [Davis1984] offer continuous-time dynamics with switching behaviors. While these models are more general in some respects, they often lack the compositional and language-theoretic semantics provided by automata-based models. Controlled SANs bridge this gap by supporting a hierarchy of semantics—from nondeterministic to stochastic—within a single automata-theoretic framework.
Controlled Automata and Language Expressiveness
The study of controlled automata and their language-theoretic properties under various policy types has a rich history in formal methods, including supervisory control theory [Ramadge1987], game-theoretic controller synthesis [Pnueli1989], and probabilistic automata [Segala1995]. Our work builds on these foundations by systematically classifying control policies (e.g., memoryless, finite-memory, stack-augmented, tape-augmented) and characterizing their expressive power in terms of accepted languages over finite and infinite words.
The hierarchy of language families induced by these policies mirrors classical distinctions in automata theory (e.g., regular, context-free, and recursively enumerable languages) and connects to foundational results on $\omega$-automata and Büchi acceptance [Thomas1990, Vardi1986Automata, Rabin1969].
These classical results provide the theoretical basis for understanding the expressive power of Controlled SANs. Depending on the class of control policy, the corresponding accepted languages may range from regular to recursively enumerable. This structured view allows a principled analysis of the interplay between control capabilities and probabilistic dynamics in both finite and infinite behaviors.
Remark 2.1.
The increasing expressiveness induced by richer classes of control policies has direct implications for the decidability and complexity of language-acceptance problems. For example, while memoryless control corresponds to regular languages with well-understood algorithmic properties, stack- or tape-augmented control can reach the level of context-free or even recursively enumerable languages, where emptiness and equivalence problems become much harder. This expressiveness–complexity trade-off forms a conceptual foundation for later undecidability and hardness results developed in this paper.
Pushdown Automata: Probabilistic and Controlled Extensions
Pushdown automata (PDAs) provide a classical foundation for modeling infinite-state systems with stack-based control. A well-studied subclass is Basic Process Algebra (BPA), which can be viewed as a special case of PDAs where the stack always contains a single nonterminal at the top and its evolution is determined by leftmost rewriting according to context-free rules in Greibach normal form.
Probabilistic BPA (pBPA) extends BPA by introducing probability distributions over rewrite rules, thereby inducing infinite-state Markov decision processes with structured stack discipline. These models have been extensively investigated in the literature and admit decidability results for several qualitative properties.
More general classes, such as stochastic BPA games [brazdil2010qualitative] and probabilistic pushdown automata (pPDA) [etessami2005recursive, etessami2009recursive, esparza2014probabilistic], capture recursion, nondeterminism, and probabilistic branching. These models correspond to infinite-state Markov decision processes with well-defined semantics and decidable reachability and $\omega$-regular properties for important subclasses.
Our Controlled Stochastic Automata (CSA) framework can be viewed as a unifying generalization of these probabilistic pushdown models, incorporating policy-based control and stochastic timing into the semantic hierarchy. This allows CSA to subsume both finite-state CTMDPs and infinite-state pushdown extensions within a single formalism.
Formal Methods in Safety-Critical Systems
Finally, formal modeling and verification of safety-critical systems has increasingly emphasized the integration of stochastic timing, uncertainty, and control [seshia2022cyberphysical, Clarke1999]. Tools like PRISM [Kwiatkowska2011], MODEST [Hahn2010], and Storm [Dehnert2017] support a subset of these features, often limited to MDP-based models. Controlled SANs provide a formalism that aligns more closely with the practical demands of such systems—offering compositionality, distributed semantics, and structured control—while retaining formal rigor and support for verification and synthesis. One of the most prominent tools for modeling and analyzing SAN-based systems is the Möbius Modeling Framework [Deavours2002Mobius], which provides a compositional, extensible environment for performance and dependability analysis. Möbius has significantly influenced the modeling community and offers a natural foundation for extensions such as Controlled Stochastic Activity Networks (CSANs).
3 Nondeterministic Models
Controlled activity networks are a generalization of activity networks [movaghar1984stochastic, movaghar2001stochastic], incorporating control mechanisms in a nondeterministic setting.
Model Structure
Throughout this paper, $\mathcal{N}$} denotes the set of natural numbers and $\mathcal{R}_{+}$} represents the set of non-negative real numbers.
Definition 3.1.
A controlled activity network is a 9-tuple
$K=(P,IA,TA,CA,IG,OG,IR,IOR,TOR)$
where:
  • $P$ is a finite set of places,
  • IAIA is a finite set of instantaneous activities,
  • TATA is a finite set of timed activities,
  • CACA is a finite set of control actions.
  • IGIG is a finite set of input gates. Each input gate has a finite number of inputs. To each $G\in IG$, with $m$ inputs, is associated a function $f_{G}:\;{\mathcal{N}}^{m}\longrightarrow{\mathcal{N}}^{m}$, called the function of $G$, and a predicate $g_{G}:\;{\mathcal{N}}^{m}\longrightarrow\{true,false\}$, called the enabling predicate of $G$,
  • OGOG is a finite set of output gates. Each output gate has a finite number of outputs. To each $G\in OG$, with $m$ outputs, is associated a function $f_{G}:\;{\mathcal{N}}^{m}\longrightarrow{\mathcal{N}}^{m}$, called the function of $G$,
  • $IR\subseteq P\times\{1,\ldots,|P|\}\times IG\times(IA\cup TA)$ is the input relation. $IR$ satisfies the following conditions:
    • for any $(P_{1},i,G,a)\in IR$ such that $G$ has $m$ inputs, $i\leq m$,
    • for any $G\in IG$ with $m$ inputs and $i\in{\mathcal{N}},$ $i\leq m$, there exist $a\in(IA\cup TA)$ and $P_{1}\in P$ such that $(P_{1},i,G,a)\in IR$,
    • for any $(P_{1},i,G_{1},a),(P_{1},j,G_{2},a)\in IR$, $i=j$ and $G_{1}=G_{2}$,
  • $IOR\subseteq IA\times OG\times\{1,\ldots|P|\}\times P$ is the instantaneous output relation.
  • $TOR\subseteq TA\times CA\times OG\times\{1,\ldots|P|\}\times P$ is the timed output relation.
    IORIOR and $TOR$ above satisfy the following conditions:
    • for any $(a,G,i,P_{1})\in IOR$ such that $G$ has $m$ outputs, $i\leq m$, item for any $(a,G,c,i,P_{1})\in TOR$ such that $G$ has $m$ outputs, $i\leq m$,
    • for any $G\in OG$ with $m$ outputs and $i\in{\mathcal{N}}$, $i\leq m$, either there exist $a\in IA$ and $P_{1}\in P$ such that $(a,G,i,P_{1})\in IOR$ or there exist $a\in TA$, $c\in CA$ and $P_{1}\in P$ such that $(a,G,c,i,P_{1})\in TOR$, but not both,
    • for any $(a,G_{1},i,P_{1}),(a,G_{2},j,P_{1})\in IOR,$ $i=j$ and $G_{1}=G_{2}$.
    • for any $(a,G_{1},c_{1},i,P_{1}),(a,G_{2},c_{2},j,P_{1})\in TOR,$ $i=j$, $c_{1}=c_{2}$ and $G_{1}=G_{2}$.
Graphically, a controlled activity network is represented as follows.
A place is depicted as . An instantaneous activity
is represented as and a timed activity as  . A control action is shown as connected to the right side of a timed activity. An input gate
with $m$ inputs is shown as
...mm11...and an output gate with
$m$ outputs as ...m1....
Except for the inclusion of control actions, the structure of controlled activity networks is the same as the structure of activity networks [movaghar2001stochastic]. As an example, consider the graphical representation of a controlled activity network as in Figure 1. $P1$, $P2$, $P3$, $P4$, $P5$, $P6$, $P7$, and $P8$ are places. $T1$, $T2$, $T3$, and $T4$ are timed activities, and $I1$, and $I2$ are instantaneous activities. $c1$, $c2$, $c3$, and $c4$ are control actions. $G1$ is an input gate with only one input, and $G2$ is an output gate with only one output. The enabling predicates, if any, and functions of these gates are indicated in a table called a “Gate Table” as depicted in Figure 1.
A directed line from a place to an activity represents a special input gate with a single input and an enabling predicate $g$ and a function $f$ such that $g(x)=true$, iff $x\geq 1$, and f(x)=x1f(x)=x-1 (e.g., directed lines from $P1$ to $I1$ and from $P2$ to $I2$). A directed line from a timed activity and a control action to a place represents a special output gate with a single output and a function $f$ such that $f(x)=x+1$ (e.g., the directed lines from $T1$ and $c1$ to $P1$ and the directed lines from $T1$ and $c2$ to $P3$). These special gates are referred to as “standard” gates.
Consider a controlled activity network as in Definition  3.1. Suppose, $(P_{k},k,G,a)\in IR$. Then, in a graphical representation, place $P_{k}$ is linked to the $k$-th input of an input gate $G$ whose output is connected to activity $a$. $P_{k}$ is said to be an input place of $a$ and $G$ is referred to as an input gate of $a$. For example, let $IR$ be the input relation of the model of Figure 1. Then, $(P4,1,G2,T2)\in IR$, $P4$, $P5$, and $P6$ are input places of $T2$, and $G2$ is an input gate of $T2$. Similarly, suppose $(a,c,G,k,P_{k})\in TOR$. Then, in a graphical representation, the timed activity $a$ via control action $c$ is connected to the input of an output gate $G$ whose $k$-th output is linked to a place $P_{k}$. $P_{k}$ is said to be an output place of the timed activity $a$, and gate $G$ is referered to as an output gate of $a$ via control action $c$. For example, in the model of Figure 1, $(T1,c1,G1,1,P2)\in TOR$, that is, a timed activity $T1$ via control action $c1$ is linked to the input of an output gate $G1$ whose only output is connected to place $P2$. $G1$ is an output gate of $T1$ via $c1$, and $P2$ is an output place of $T1$ via $c1$. Note also that $(T1,c1,G^{\prime},1,P1),(T1,c2,G^{\prime\prime},1,P3)\in TOR$, where $G^{\prime}$ and $G^{\prime\prime}$ are standard output gates, which are the output gates of $T1$ via $c1$ and $c2$, respectively, and that $P1$ and $P3$ are the output places of $T1$ via $c1$ and $c2$, respectively.
Like Petri nets, we have a notion of “marking” for controlled activity networks.
Definition 3.2.
Consider a controlled activity network as in Definition  3.1. A marking is a function
$\mu:\;P\longrightarrow{\mathcal{N}}.$
It is often convenient to characterize a marking $\mu$ as a vector, that is, $\mu=(\mu_{1},\ldots,\mu_{n})$, where $\mu_{i}=\mu(P_{i})$, $i=1,\ldots,n$, and $P=\{P_{1},\ldots,P_{n}\}$. In a graphical representation, a marking is characterized by tokens (dots) inside places. The number of tokens in a place represents the marking of that place (e.g., marking $(0,,0,0,1,1,1,0,0)$ in Figure 1). An activity is “enabled” in a marking if the enabling predicates of its input gates are true in that marking.
More formally, we have:
Definition 3.3.
Consider a controlled activity network as in Definition  3.1. $a\in(IA\cup TA)$ is enabled in a marking $\mu$ if for any input gate $G$ of $a$ with $m$ inputs and an enabling predicate $g_{G}$,
$g_{G}(\mu_{1}.\dots,\mu_{m})=true,$
where $\mu_{k}=\mu(P_{k})$, for some $P_{k}\in P$ such that $(P_{k},k,G,a)\in IR$, $k=1,\dots,m$.
An activity is disabled in a marking if it is not enabled in that marking. A marking is stable if no instantaneous activity is enabled in that marking. A marking is unstable if it is not stable. For example, $(0,,0,1,0,1,1,0,0)$ is an unstable marking but $(0,,0,0,1,1,1,0,0)$ is a stable marking in the model of Figure 1. In the former marking, timed activity $T1$ and instantaneous activity $I3$ are enabled, but in the latter marking, timed activities $T1$ and $T2$ are enabled.
Model Behavior
A controlled activity network with a marking is a dynamic system. A marking changes only if an activity completes. In a stable marking, only one of the enabled timed activities is allowed to complete. When there is more than one enabled timed activity, the choice of which activity to complete first is done nondeterministically. In an unstable marking, only one of the enabled instantaneous activities may complete (i.e., enabled instantaneous activities have priority over enabled timed activities for completion). When there is more than one enabled instantaneous activity, the choice of which activity to complete first is also done nondeterministically. When an activity completes, it may change the marking of its input and output places. This change is governed by the functions of its input gates and output gates, and is done in two steps as follows. First, the marking of its input places may change due to the functions of its input gates, resulting in an intermediary marking. Next, in this latter marking, the marking of its output places may also change due to the functions of its output gates, resulting in a final marking after the completion of that activity. More specifically, let us consider a controlled activity network as in Definition  3.1. Suppose an activity $a$ completes in a marking $\mu$. The next marking $\mu^{\prime}$ is determined in two steps as follows. First, an intermediary marking $\mu^{\prime\prime}$ is obtained from $\mu$ by the functions of the input gates of $a$. Then, $\mu^{\prime}$ is determined depending on whether $a$ is an instantaneous activity or a timed activity. In the first case, $\mu^{\prime}$ will be determined from $\mu^{\prime\prime}$ by the functions of the output gates of $a$. In the latter case, first, a control action is chosen. When there is more than one control action, the choice is made according to a given policy (to be defined later in this section). $\mu^{\prime}$ will then be determined from $\mu^{\prime\prime}$ by the functions of the output gates of $a$ via the chosen control action.
More formally, $\mu^{\prime\prime}$ and $\mu^{\prime}$ are defined as follows:
  • for any $P_{1}\in P$ which is not an input or output place of $a$,
    $\mu^{\prime\prime}(P_{1})=\mu^{\prime}(P_{1})=\mu(P_{1}),$
  • for any input gate $G$ of $a$ with $m$ inputs and a function $f_{G}$,
    $f_{G}(\mu_{1},\ldots,\mu_{m})=(\mu^{\prime\prime}_{1},\ldots,\mu^{\prime\prime}_{m}),$
    where $\mu_{k}=\mu(P_{k})$ and $\mu^{\prime\prime}_{k}=\mu^{\prime\prime}(P_{k})$ such that $(P_{k},k,G,a)\in IR$, $k=1,\ldots,m$,
  • for any output gate $G$ of an instatnaneous activity $a$ with $m$ outputs and a function $f_{G}$,
    $f_{G}(\mu^{\prime\prime}_{1},\dots,\mu^{\prime\prime}_{m})=(\mu^{\prime}_{1},\ldots,\mu^{\prime}_{m}),$
    where $\mu^{\prime\prime}_{k}=\mu^{\prime\prime}(P_{k})$ and $\mu^{\prime}_{k}=\mu^{\prime}(P_{k})$ such that $(a,G,k,P_{k})\in IOR$, $k=1,\ldots,m$.
  • for any output gate $G$ of a timed activity $a$ choosing control action $c$ with $m$ outputs and a function $f_{G}$,
    $f_{G}(\mu^{\prime\prime}_{1},\dots,\mu^{\prime\prime}_{m})=(\mu^{\prime}_{1},\ldots,\mu^{\prime}_{m}),$
    where $\mu^{\prime\prime}_{k}=\mu^{\prime\prime}(P_{k})$ and $\mu^{\prime}_{k}=\mu^{\prime}(P_{k})$ such that $(a,c,G,k,P_{k})\in TOR$, $k=1,\ldots,m$.
The above summarizes the behavior of a controlled activity network. As an example, consider the model of Figure 1 with a marking $(0,0,0,,1,1,1,0,0)$. In this marking, the timed activities $T1$ and $T2$ are enabled. Any of these activities may be completed. Suppose $T1$ completes first. Then, one of the control actions $c1$ and $c2$ is chosen. Let us assume $c1$ is chosen. Then, $P1$ and $P2$ each gain a token, and the marking changes to marking $(1,1,0,1,1,1,0,0)$, as in Figure 2, which is an unstable marking. In this marking, timed activities $T1$ and $T2$ and instantaneous activities $I1$ and $I2$ are enabled. Only $I1$ or $I2$ is allowed to be completed next. Suppose $I2$ completes next. The resulting marking will be $(1,0,0,1,2,1,0,0)$, as in Figure 3, which is an unstable marking. In this marking, only $I1$ is allowed to complete, resulting in the stable marking $(0,0,0,2,2,1,0,0)$, as in Figure 4.
Following the above discussion, a marking $\mu^{\prime}$ is said to be reachable from a marking $\mu$ under a string of activities $a_{1}\dots a_{n}$, if the successive completion of $a_{1},\ldots,a_{n}$ changes the marking of the network from $\mu$ to $\mu^{\prime}$ under some possible choice of a sequence of control actions. $\mu^{\prime}$ is said to be reachable from $\mu$ if $\mu^{\prime}$ is reachable from $\mu$ under a string of activities or $\mu^{\prime}=\mu$. For example, $(0,0,0,2,2,1,0,0)$ is reachable from $(0,0,0,1,1,1,0,0)$ under $T1I2I1$, and $(1,0,0,1,2,1,0,0)$ is reachable from $(0,0,0,,1,1,1,0,0)$.
Semantic Models
The behavior of a controlled activity network is concerned with how various stable markings are reached from each other due to the completion of timed activities. To study this behavior more formally, we take advantage of the notion of a controlled automaton.
Definition 3.4.
A controlled automaton is a 5-tuple $S=(Q,A,C,\rightarrow,Q_{0})$ where:
  • $Q$ is the set of states,
  • AA is the activity alphabet,
  • $C$ is the control action alphabet,
  • $\rightarrow\subseteq Q\times A\times C\times Q$ is the transition relation,
  • $Q_{0}\subseteq Q$ is the set of initial states.
We use the term ‘automaton’ in a general sense to describe state-transition structures governed by activity completions and control actions, without requiring acceptance criteria. We also write $q\xrightarrow{a,c}q^{\prime}$ to denote $(q,a,c,q^{\prime})\in\rightarrow$. In this case, we say that state $q^{\prime}$ is immediately reachable from a state $q$ under an activity $a$ and control action $c$, and that activity $a$ is enabled in state $q$. For $q_{i}\in Q$, $a_{i}\in A$ and $c_{i}\in C$, $0\leq i\leq n$, $q_{0}\xrightarrow{a_{0},c_{0}}q_{1}\xrightarrow{a_{1},c_{1}}\ldots q_{n-1}\xrightarrow{a_{n-1},c_{n-1}}q_{n}$ is called a finite execution in $S$ with lenght $n$ and $(q_{0},a_{0})(q_{1},a_{1})\ldots(q_{n-1},a_{n-1})q_{n}$ is referred to as a finite run of this execution with length $n$ in $S$. The above execution may also be denoted simply as $(q_{0},a_{0},c_{0})(q_{1},a_{1},c_{1})\ldots(q_{n-1},a_{n-1},c_{n-1})q_{n}$.
Bisimulation
Next, we define a notion of equivalence for controlled automata based on the concept of bisimulation, which is similar to the one proposed for automata [Milner1989, baier2008principles].
Definition 3.5.
Let $S=(Q,A,C,\rightarrow,Q_{0})$ and $S^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},\rightarrow^{\prime},Q^{\prime}_{0})$ be two controlled automata with the same activity alphabet and set of control action alphabet(i.e., $A=A^{\prime}$ and $C=C^{\prime}$). $S$ and $S^{\prime}$ are said to be equivalent if there exists a symmetric binary relation $\gamma$ on $Q\cup Q^{\prime}$ such that:
  • for any $q\in Q$, there exists a $q^{\prime}\in Q^{\prime}$ such that $(q,q^{\prime})\in\gamma$; also, for any $q^{\prime}\in Q^{\prime}$, there exists a $q\in Q$ such that $(q^{\prime},q)\in\gamma$,
  • for any $q\in Q_{0}$, there exists a $q^{\prime}\in Q^{\prime}_{0}$ such that $(q,q^{\prime})\in\gamma$; also, for any $q^{\prime}\in Q^{\prime}_{0}$, there exists a $q\in Q_{0}$ such that $(q^{\prime},q)\in\gamma$,
  • for any $q_{1},q_{2}\in Q$, $q^{\prime}_{1}\in Q^{\prime}$, $a\in A$ and $c\in C$ such that $(q_{1},q^{\prime}_{1})\in\gamma$ and $(q_{1},a,c,q_{2})\in\rightarrow$, there exists $q^{\prime}_{2}\in Q^{\prime}$ such that $(q_{2},q^{\prime}_{2})\in\gamma$ and $(q^{\prime}_{1},a,c,q^{\prime}_{2})\in\rightarrow^{\prime}$; also, for any $q^{\prime}_{1},q^{\prime}_{2}\in Q^{\prime}$, $q_{1}\in Q$, $a\in A$ and $c\in C$ such that $(q^{\prime}_{1},q_{1})\in\gamma$ and $q^{\prime}_{1},a,c,q^{\prime}_{2})\in\rightarrow^{\prime}$, there exists $q_{2}\in Q$ such that $(q^{\prime}_{2},q_{2})\in\gamma$ and $(q_{1},a,c,q_{2})\in\rightarrow$.
$\gamma$ above is said to be a bisimulation between the two controlled automata $S$ and $S^{\prime}$ above. $S$ and $S^{\prime}$ are isomorphic if $\gamma$ is a bijection.
Proposition 3.6.
Let $\mathcal{E_{S}}$ denote a relation on the set of all controlled automata such that $(S_{1},S_{2})\in\mathcal{E_{S}}$ if and only if $S_{1}$ and $S_{2}$ are equivalent controlled automata in the sense of Definition 3.5. Then $\mathcal{E_{S}}$ will be an equivalence relation.
Proposition 3.7.
Let $S=(Q,A,C,\rightarrow,Q_{0})$ and $S^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},\rightarrow^{\prime},Q^{\prime}_{0})$ be two equivalent controlled automata with the same activity alphabet and set of control actions (i.e., $A=A^{\prime}$ and $C=C^{\prime}$) under a bisimulation $\gamma$. Then, we have:
  • for any finite execution $(q_{0},a_{0},c_{0})(q_{1},a_{1},c_{1})\ldots(q_{n-1,}a_{n-1},c_{n-1})q_{n}$ in $S$, there exists a finite execution $(q^{\prime}_{0},a_{0},c_{0})(q^{\prime}_{1},a_{1},c_{1})\ldots(q^{\prime}_{n-1},a_{n-1},c_{n-1})q^{\prime}_{n}$ in $S^{\prime}$ such that $(q_{i},q^{\prime}_{i})\in\gamma$, $0\leq i\leq n$, and
  • for any finite execution $(q^{\prime}_{0},a_{0},c_{0})(q^{\prime}_{1},a_{1},c_{1})\ldots(q^{\prime}_{n-1},a_{n-1},c_{n-1})q^{\prime}_{n}$ in $S^{\prime}$, there exists a finite execution $(q_{0},a_{0},c_{0})(q_{1},a_{1},c_{1})\ldots(q_{n-1},a_{n-1},c_{n-1})q_{n}$ in $S$ such that $(q^{\prime}_{i},q_{i})\in\gamma$, $0\leq i\leq n$.
Operational Semantics
We are now in a position to formalize the notion of the behavior of a controlled activity network as follows.
Definition 3.8.
Let $(K,\mu_{0})$ denote a controlled activity network $K$ with an initial marking $\mu_{0}$ where $K$ is defined as in Definition  3.1. $(K,\mu_{0})$ is said to realize a controlled automaton $S=(Q,A,C,\rightarrow,Q_{0})$ where:
  • $Q$ is the set of all stable markings of $K$ which are reachable from $\mu_{0}$ and a state $\Delta$ if, in $K$, an infinite sequence of instantaneous activities can be completed in a marking reachable from $\mu_{0}$,
  • AA is the set of timed activities of $K$,
  • $C$ is the set of control actions of $K$,
  • for any $\mu,\mu^{\prime}\in Q$ , $a\in A$ and $c\in C$, $(\mu,a,c,\mu^{\prime})\in,\rightarrow$, iff, in $K$, $\mu^{\prime}$ is reachable from $\mu$ under a string of activities $ax$, where $x$ is a (possibly an empty) string of instantaneous activities; $(\mu,a,c,\Delta)\in\rightarrow$, iff, in $K$, a sequence of activities $ay$ can complete in $\mu$, where $y$ is an infinite sequence of instantaneous activities,
  • $Q_{0}$ is the set of all stable markings of $K$ which are reachable from $\mu_{0}$ under a (possibly an empty) string of instantaneous activities and $\Delta$ if, in $K$, an infinite sequence of instantaneous activities can complete in $\mu_{0}$.
The above definition implies a notion of equivalence for controlled activity networks as follows.
Definition 3.9.
Two controlled activity networks are equivalent if they realize equivalent controlled automata.
Proposition 3.10.
Let $\mathcal{E_{K}}$ denote a relation on the set of all controlled activity networks such that $(K_{1},K_{2})\in\mathcal{E_{K}}$ if and only if $K_{1}$ and $K_{2}$ are equivalent controlled activity networks in the sense of Definition 3.9. Then $\mathcal{E_{K}}$ will be an equivalence relation.
Modeling Power
The following concepts help specify the modeling power of controlled activity networks.
Definition 3.11.
A controlled automaton is said to be computable if it has a computable transition relation and an enumerable set of initial states.
Definition 3.12.
A controlled activity network is said to be computable if the enabling predicates and functions of all of its input gates and the functions of all of its output gates are computable.
The following theorem more precisely characterizes the modeling power of controlled activity networks
Theorem 3.13.
Any computable controlled automaton is isomorphic to a controlled automaton realized by a computable controlled activity network with some initial marking.
Proof Sketch.
The proof relies on a notable result about Petri nets—specifically, that the class of extended Petri nets [Peterson1981], which includes inhibitor arcs, can simulate nondeterministic Turing machines [Hack1975, Sipser2013]. Consequently, an activity network that incorporates instantaneous activities, standard gates, and inhibitor gates can likewise simulate a nondeterministic Turing machine and, therefore, represent any computable relation. This result is instrumental in the proof of the theorem. For the full proof, see Appendix A. ∎
The above proof shows that, in terms of behavioral equivalence, any system modeled by a computable controlled activity network can also be represented using a model that includes only standard gates and inhibitor gates. However, achieving this representation may require a significantly larger and more complex model, potentially at the cost of modeling clarity and convenience. Formally, this leads to the following corollary:
Corollary 3.14.
Any computable controlled activity network is isomorphic to a controlled activity network that has only standard and inhibitor gates.
Policy Types
Next, we consider the notion of a policy and policy types in a controlled automaton as follows.
Definition 3.15.
Let $S=(Q,A,C,\rightarrow,Q_{0})$ be a controlled automaton. A (history-dependent) policy $\pi$ is defined as
$\pi:(Q\times A)^{*}\rightarrow C.$
For $q_{i}\in Q$, $a_{i}\in A$ and $c_{i}\in C$, $0\leq i\leq n$, an execution $q_{0}\xrightarrow{a_{0},c_{0}}q_{1}\xrightarrow{a_{1},c_{1}}\ldots q_{n-1}\xrightarrow{a_{n-1},c_{n-1}}q_{n}$ in $S$ is defined under a policy $\pi$ if $c_{i-1}=\pi(h_{i})$, where $h_{i}=(q_{0},a_{0})\ldots(q_{i-1},a_{i-1})$ for all $0<i\leq n$. $\pi$ is called memoryless if $\pi:Q\rightarrow C$.
Definition 3.16.
Let $S=(Q,A,C,\to,Q_{0})$ and $S^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},\to^{\prime},Q_{0}^{\prime})$ be equivalent controlled automata sharing the same activity alphabet and control set, i.e. $A=A^{\prime}$ and $C=C^{\prime}$. Let $\gamma\subseteq Q\times Q^{\prime}$ be a bisimulation between $S$ and $S^{\prime}$. For the sake of brevity, we denote a finite execution $(q_{0},a_{0},c_{0})(q_{1},a_{1},c_{1})\dots(q_{n-1},a_{n-1},c_{n-1})q_{n}$ of $S$ as $(q_{i},a_{i},c_{i})_{i=0}^{n-1}\,q_{n}$, where for each $0\leq i<n$ we have $(q_{i},a_{i},c_{i},q_{i+1})\in\to$. Recall the history available at step $i$ for this execution is defined as
$h_{i}\;=\;(q_{0},a_{0})\,(q_{1},a_{1})\cdots(q_{i-1},a_{i-1})\quad(\text{with }h_{0}=\varepsilon),$
and analogously $h^{\prime}_{i}$ for an execution of $S^{\prime}$.
Two (history-dependent) policies $\pi$ on $S$ and $\pi^{\prime}$ on $S^{\prime}$ are said to be equivalent under $\gamma$ if for every pair of executions
$(q_{i},a_{i},c_{i})_{i=0}^{n-1}\,q_{n}\text{ of }S\quad\text{and}\quad(q^{\prime}_{i},a_{i},c_{i})_{i=0}^{n-1}\,q^{\prime}_{n}\text{ of }S^{\prime}$
of the same length and with $(q_{i},q^{\prime}_{i})\in\gamma$ for all $0\leq i\leq n$, the recommended controls coincide at every step:
$\pi(h_{i})\;=\;\pi^{\prime}(h^{\prime}_{i})\;=\;c_{i-1}\qquad\text{for all }\,0<i<n.$
We now define three types of policies for controlled automata by augmenting these models with finite-state automata, pushdown automata, and Turing machines [Sipser2013] as follows.
Definition 3.17.
A finite-memory policy is a tuple $\pi=(M,m_{0},\delta,\gamma)$ where:
  • $M$: finite set of memory states,
  • $m_{0}\in M$: initial memory state,
  • $\delta:M\times(Q\times A)\rightarrow M$: memory update function,
  • $\gamma:Q\times M\rightarrow C$: control output function.
Definition 3.18.
A stack-augmented policy is a tuple $\pi=(\Gamma,\bot,m_{0},\delta,\gamma)$ where:
  • $\Gamma$: finite stack alphabet with bottom symbol $\bot$,
  • $m_{0}$: initial configuration (state and stack),
  • $\delta:\Gamma\times Q\times A\rightarrow\Gamma^{*}$: stack update function depending only on the top stack symbol,
  • $\gamma:Q\times\Gamma\rightarrow C$: control output function depending only on the top stack symbol.
Definition 3.19.
A tape-augmented policy is a tuple $\pi=(\Gamma,m_{0},\delta,\gamma)$ where:
  • $\Gamma$: finite tape alphabet,
  • $m_{0}$: initial tape configuration and head position,
  • $\delta:\Gamma\times Q\times A\rightarrow\Gamma\times\{L,R\}$: tape update and head move function depending only on the scanned symbol,
  • $\gamma:Q\times\Gamma\rightarrow C$: control output function depending only on the current scanned tape symbol.
Nondeterministic Languages on Finite Words
In the remainder of this section, to further explore formal language recognition by controlled automata, we assume that the automaton has a finite set of states along with a designated subset $F$, referred to as the set of accepting states.
Definition 3.20.
Let $S=(Q,A,C,P,Q_{0},F)$ be a controlled automaton with a set of accepting states ($F\subseteq Q$). Given a policy $\pi$ in $S$, the language accepted by $S$ under $\pi$ is defined as:
$L_{\pi}(S)=\bigl\{\,a_{0}a_{1}\ldots a_{n-1}\in A^{*}\ \bigm|$  $\exists\text{ a run }(q_{0},a_{0})\ldots(q_{n-1},a_{n-1})\,q_{n}\text{ in }S\text{ under }\pi$
$\text{ such that }q_{0}\in Q_{0}\text{ and }\ q_{n}\in F\,\bigr\}.$
Definition 3.21.
A controlled automaton $S=(Q,A,C,\rightarrow,q_{0})$ is deterministic if for every state $q\in Q$, input symbol $a\in A$, and control action $c\in C$, there exists at most one successor state $q^{\prime}\in Q$ such that $(q,a,c,q^{\prime})\in\rightarrow$.
Equivalently, the transition relation $\rightarrow$ defines a partial function
$\delta:Q\times A\times C\rightharpoonup Q.$
If $\delta$ is total, the automaton is said to be completely deterministic.
Proposition 3.22.
Let $S$ be deterministic and $\pi:(Q\times A)^{*}\rightarrow C$. Then for every initial state $q_{0}\in Q_{0}$ and string $a_{0}a_{1}\ldots a_{n-1}\in A^{*}$, there is at most one execution:
$q_{0}\xrightarrow{a_{0},\pi(h_{1})}q_{1}\xrightarrow{a_{1},\pi(h_{2})}\ldots\xrightarrow{a_{n-1},\pi(h_{n})}q_{n},$
where $h_{i}=(q_{0},a_{0})\ldots,(q_{i-1}a_{i-1})$ for $0<i\leq n$.
Throughout this paper, $\subset$ denotes proper inclusion, i.e., $A\subset B$ iff $A\subseteq B$ and $A\neq B$. The following properties hold:
Proposition 3.23.
Let $S$ be a controlled automaton. Define the following language classes:
$L_{0}(S)$ $=\{L_{\pi}(S)\mid\pi\text{ is a memoryless policy}\},$
$L_{F}(S)$ $=\{L_{\pi}(S)\mid\pi\text{ is a finite-memory policy}\},$
$L_{\text{Stack}}(S)$ $=\{L_{\pi}(S)\mid\pi\text{ is a stack-augmented policy}\},$
$L_{\text{Tape}}(S)$ $=\{L_{\pi}(S)\mid\pi\text{ is a tape-augmented policy}\},$
$L_{H}(S)$ $=\{L_{\pi}(S)\mid\pi\text{ is a history-dependent policy}\}.$
Then we have:
$L_{0}(S)\subset L_{F}(S)\subset L_{\text{Stack}}(S)\subset L_{\text{Tape}}(S)\subset L_{H}(S)$
Proof Sketch..
This result follows directly from the arguments in the proof sketch of Theorem 3.29. ∎
Definition 3.24.
Let $S=(Q,A,C,\rightarrow,Q_{0},F)$ and $S^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},\rightarrow^{\prime},Q^{\prime}_{0},F^{\prime})$ be two controlled automata with accepting states $F$ and $F^{\prime}$, respectively, and the same activity alphabet and set of control actions (i.e., $A=A^{\prime}$ and $C=C^{\prime}$). Let $(Q,A,C,\rightarrow,Q_{0})$ and $(Q^{\prime},A^{\prime},C^{\prime},\rightarrow^{\prime},Q^{\prime}_{0})$ be two equivalent controlled automata under a bisimulation $\gamma$ as in Definition  3.5. Then, $S$ and $S^{\prime}$ are said to be equivalent if:
  • for any $q\in F$, there exists $q^{\prime}\in F^{\prime}$ where $(q,q^{\prime})\in\gamma$; also for any $q^{\prime}\in F^{\prime}$, there exists $q\in F$ where $(q^{\prime\prime},q)\in\gamma$.
Proposition 3.25.
Let $S$ and $S^{\prime}$ be two equivalent controlled automata with accepting states as in Definition 3.24. Consider the classes of languages as defined in Proposition 3.23. Then, we have:
$L_{0}(S)$ $=L_{0}(S^{\prime})$
$L_{F}(S)$ $=L_{F}(S^{\prime})$
$L_{\text{Stack}}(S)$ $=L_{\text{Stack}}(S^{\prime})$
$L_{\text{Tape}}(S)$ $=L_{\text{Tape}}(S^{\prime})$
$L_{H}(S)$ $=L_{H}(S^{\prime})$
Also, for equivalent polices $\pi$ and $\pi^{\prime}$ of $S$ and $S^{\prime}$, respectively, as in Definition 3.16, we have
$L_{0}^{\pi}(S)$ $=L_{0}^{\pi^{\prime}}(S^{\prime})$
$L_{F}^{\pi}(S)$ $=L_{F}^{\pi^{\prime}}(S^{\prime})$
$L_{\text{Stack}}^{\pi}(S)$ $=L_{\text{Stack}}^{\pi^{\prime}}(S^{\prime})$
$L_{\text{Tape}}(S)$ $=L_{\text{Tape}}^{\pi^{\prime}}(S^{\prime})$
$L_{H}^{\pi}(S)$ $=L_{H}^{\pi^{\prime}}(S^{\prime})$
Proposition 3.26.
Let $L_{H}^{\mathrm{comp}}(S)$ denote the class of languages accepted a controlled automaton $S=(Q,A,C,\rightarrow,Q_{0})$ under computable history-dependent policies $\pi:(Q\times A)^{*}\to C$. Then any such policy can be simulated by a tape-augmented policy. Therefore,
$L_{\text{Tape}}(S)=L_{H}^{\mathrm{comp}}(S).$
Global Language Families for Controlled Automata on Finite Words
We now explore the structure and hierarchy of language families defined by classes of control policies in controlled automata. We define, interpret, and compare these families, with an emphasis on the expressive power granted by different forms of memory and history-awareness in controllers. More specifically, we define global language families $\mathbb{L}_{\Pi}$ for several classes of policies $\Pi$, each corresponding to increasing expressive power in controller design.
Definition 3.27.
Let $\mathcal{S}$ denote the set of all controlled automata. For each policy class $\Pi\in\{0,F,\text{Stack},\text{Tape},H\}$, define:
$\mathbb{L}_{\Pi}:=\{L_{\Pi}(S)\mid S\in\mathcal{S}\}\subseteq\mathcal{P}(2^{A^{*}})$
where $A$ is the set of activities and each $L_{\Pi}(S)=\{L_{\pi}(S)\mid\pi\in\Pi\}$ is the class of languages accepted by controlled automaton $S$ under policies of type $\Pi$.
Interpretation
Each element of $\mathbb{L}_{\Pi}$ is a set of languages induced by a specific automaton $S$ operating under a family of policies $\Pi$. Thus, $\mathbb{L}_{\Pi}$ is a set of language families. While the union of all such languages may correspond to a classical complexity class (e.g., $\bigcup\mathbb{L}_{F}=\textbf{REG}$), we emphasize:
$\mathbb{L}_{\Pi}$ consists of language families—not individual languages.
We now characterize the union over all language families within each global class $\mathbb{L}_{\Pi}$.
Proposition 3.28.
The union of all languages accepted by controlled automata under memory-augmented policies equals the class of regular languages:
$\bigcup\mathbb{L}_{F}=\textbf{REG}$
The union of all languages accepted by controlled automata under stack-augmented policies equals the class of context-free languages:
$\bigcup\mathbb{L}_{\text{Stack}}=\textbf{CFL}$
The union of all languages accepted by controlled automata under tape-augmented policies equals the class of recursively enumerable languages:
$\bigcup\mathbb{L}_{\text{Tape}}=\bigcup\mathbb{L}_{H}^{\mathrm{comp}}=\textbf{RE}$
These results reflect the alignment between policy memory models and classical language classes: stack memory corresponds to context-free expressiveness, while tape memory corresponds to Turing-recognizable power.
Hierarchy
There is a strict hierarchy of global language families:
Theorem 3.29.
Consider the language classes as defined in Definition 3.27 above. Then, we have:
$\mathbb{L}_{0}\subset\mathbb{L}_{F}\subset\mathbb{L}_{\text{Stack}}\subset\mathbb{L}_{\text{Tape}}=\mathbb{L}_{H}^{\mathrm{comp}}\subset\mathbb{L}_{H}$
Proof Sketch..
We outline why each inclusion holds and why it is strict.
1. $\mathbb{L}_{0}\subset\mathbb{L}_{F}$: Every memoryless control policy can be regarded as a degenerate finite-memory controller with a single internal state. However, finite-memory controllers can encode bounded protocol phases that memoryless ones cannot distinguish. Strictness follows from Lemma 3.30.
2. $\mathbb{L}_{F}\subset\mathbb{L}_{\text{Stack}}$: Finite-memory control is subsumed by stack-based control, since a pushdown controller can simulate any finite-memory controller and further recognize non-regular patterns such as $\{a^{n}b^{n}\mid n\geq 0\}$. Hence, the inclusion is strict.
3. $\mathbb{L}_{\text{Stack}}\subset\mathbb{L}_{\text{Tape}}$: A Turing-power controller can simulate every pushdown controller, but not vice versa; languages such as $\{a^{n}b^{n}c^{n}\mid n\geq 0\}$ separate the two classes.
4. $\mathbb{L}_{\text{Tape}}=\mathbb{L}_{H}^{\mathrm{comp}}\subset\mathbb{L}_{H}$: Computationally augmented (Turing-power) controllers correspond exactly to effectively realizable history-dependent policies; they form a subset of the most general history-dependent ones, which may be non-computable.
Thus,
$\mathbb{L}_{0}\subset\mathbb{L}_{F}\subset\mathbb{L}_{\text{Stack}}\subset\mathbb{L}_{\text{Tape}}=\mathbb{L}_{H}^{\mathrm{comp}}\subset\mathbb{L}_{H}$
Lemma 3.30.
There exists a controlled automaton $S$ such that $L_{F}(S)$ contains a countably infinite family of distinct regular languages.
Proof Sketch..
Fix the one-letter alphabet $\Sigma=\{a\}$. For each integer $k\geq 1$, define the regular language
$L_{k}\;=\;(a^{k})^{*}\;=\;\{\,(a^{k})^{m}\mid m\geq 0\,\},$
i.e., the set of all finite words whose length is a multiple of $k$. These languages are pairwise distinct (for example, $a^{2}\in L_{2}$ but $a^{2}\notin L_{3}$), and the family $\{L_{k}\mid k\geq 1\}$ is countably infinite.
Now construct a controlled automaton $S$ that can repeatedly emit the symbol $a$ and, at any step, may either continue or stop (i.e., terminate the word). Intuitively, $S$ offers two types of actions at each step: “emit aa” or “terminate.”
For each $k\geq 1$, define a finite-memory control policy $\pi_{k}$ whose internal memory is a $k$-state counter cycling through $\{0,1,\dots,k-1\}$ modulo $k$. Under $\pi_{k}$, the controller: (i) always allows the “emit aa” action, which increments the counter modulo $k$, and (ii) permits the “terminate” action only when the counter is 0 (i.e., after emitting a multiple of $k$ symbols since the start). Therefore, under $\pi_{k}$, the set of all possible terminated traces of $S$ is exactly $L_{k}=(a^{k})^{*}$.
Because each $\pi_{k}$ uses only finitely many internal memory states, $\pi_{k}$ is a finite-memory policy; hence $L_{k}\in L_{F}(S)$ for every $k$. Since the languages $L_{k}$ are all regular, are pairwise distinct, and range over all $k\geq 1$, it follows that $L_{F}(S)$ contains a countably infinite family of distinct regular languages. ∎
Proposition 3.31.
There exists a countably infinite family of pairwise distinct regular languages that is not equal to $L_{F}(S)$ for any controlled automaton $S$.
Proof Sketch..
For each integer $k\geq 1$, let
$L_{k}\;=\;(a^{k})^{*}\;=\;\{(a^{k})^{m}\mid m\geq 0\},$
the set of all finite words over $\{a\}$ whose length is a multiple of $k$. Each $L_{k}$ is regular, and the languages $L_{k}$ are pairwise distinct (e.g. $a^{2}\in L_{2}$ but $a^{2}\notin L_{3}$). In particular, the subfamily
$\mathcal{F}\;=\;\{\,L_{p}\mid p\text{ prime}\,\}\;=\;\{\,(a^{p})^{*}\mid p\text{ prime}\,\}$
is a countably infinite family of distinct regular languages.
Assume, toward contradiction, that there is a controlled automaton $S$ such that $L_{F}(S)=\mathcal{F}$. By definition, $L_{F}(S)$ is the collection of all languages realizable by $S$ under some finite-memory control policy. In particular, for each prime $p$, there is some finite-memory policy $\pi_{p}$ under which $S$ generates exactly $L_{p}=(a^{p})^{*}$.
Now observe that a finite-memory controller that enforces $L_{p}$ necessarily counts steps modulo $p$ and only allows termination after emitting a multiple of $p$ copies of $a$. From such a controller, we can build another finite-memory controller that repeats the counting cycle $p$ times before allowing termination. This modified controller permits termination only after emitting a multiple of $p^{2}$ copies of $a$, and therefore enforces exactly $L_{p^{2}}=(a^{p^{2}})^{*}$.
Since this modified controller still has only finitely many internal states, it is also a finite-memory control policy for $S$. Hence $L_{p^{2}}\in L_{F}(S)$ for each prime $p$. But $p^{2}$ is not prime, so $L_{p^{2}}\notin\mathcal{F}$ by construction. This contradicts our assumption that $L_{F}(S)=\mathcal{F}$.
Therefore, no single controlled automaton $S$ can satisfy $L_{F}(S)=\mathcal{F}$. We conclude that there exists a countably infinite family of pairwise distinct regular languages (namely, $\mathcal{F}$) that is not equal to $L_{F}(S)$ for any $S$. ∎
We next formalize the closure properties of language families induced by controlled automata and classes of control policies. Closure under standard language operations such as union, intersection, complement, and concatenation provides a natural means of comparing expressive power across policy classes. To capture both model-level and language-level perspectives, we distinguish between two related notions of closure. The first, called uniform closure, requires that the same constructive transformation of plants realizes a given operation across all admissible languages. The second, called standard or pointwise closure, coincides with the classical definition from automata and language theory, requiring only that the resulting language belongs to the same global family. Both notions are presented below in a unified framework.
Closure Properties of Controlled Language Families
Having characterized the expressive hierarchy of controlled automata, we now examine the algebraic behavior of the induced language families under standard language operations. Closure properties provide a compositional view of how control capabilities interact with the structural transformations of languages such as union, intersection, concatenation, and complement.
Definition 3.32 (Uniform and Standard Closure under Language Operations).
Let $\mathcal{S}$ denote the class of controlled automata (plants), and for each policy class $\Pi$ let $L_{\Pi}(S)$ be the family of languages realizable by policies $\pi\in\Pi$ on a plant $S\in\mathcal{S}$.
(a) Uniform Closure. The family $\mathbb{L}_{\Pi}=\{L_{\Pi}(S)\mid S\in\mathcal{S}\}$ is uniformly closed under a $k$-ary language operation $Op:(P(A^{*}))^{k}\!\to\!P(A^{*})$ if there exists a computable constructor
$\Phi_{Op}:\mathcal{S}^{k}\longrightarrow\mathcal{S}$
such that for all plants $S_{1},\dots,S_{k}\in\mathcal{S}$ and for all $L_{i}\in L_{\Pi}(S_{i})$ ($1\!\leq\!i\!\leq\!k$),
$Op(L_{1},\dots,L_{k})\in L_{\Pi}\bigl(\Phi_{Op}(S_{1},\dots,S_{k})\bigr).$
Uniform closure thus requires a constructive correspondence between the plant structures: the same syntactic transformation $\Phi_{Op}$ must realize the operation $Op$ for all admissible languages.
(b) Standard (Pointwise) Closure. The global language family
$\mathcal{C}_{\Pi}=\bigcup_{S\in\mathcal{S}}L_{\Pi}(S)$
is closed under $Op$ if for all $L_{1},\dots,L_{k}\in\mathcal{C}_{\Pi}$,
$Op(L_{1},\dots,L_{k})\in\mathcal{C}_{\Pi}.$
Equivalently, for every choice of plants $S_{1},\dots,S_{k}\in\mathcal{S}$ and policies $\pi_{1},\dots,\pi_{k}\in\Pi$, there exist a plant $S^{\prime}\in\mathcal{S}$ and a policy $\pi^{\prime}\in\Pi$ such that
$Op\!\bigl(L_{\pi_{1}}(S_{1}),\dots,L_{\pi_{k}}(S_{k})\bigr)=L_{\pi^{\prime}}(S^{\prime}).$
Unlike the uniform case, no structural relation between $S^{\prime}$ and the original plants are required.
Remark 3.33.
The distinction between standard and uniform closure reflects two complementary levels of compositional reasoning within the finite-word setting. Standard closure is a language-level property: it asserts that the family of all realizable languages under a given policy class $\Pi$ is closed under a specified language operation, independently of how the underlying plant structures are related. Uniform closure, by contrast, is a stronger model-level property: it requires the existence of a single constructive transformation $\Phi_{\mathrm{Op}}$ on plants that systematically realizes the effect of the operation across all admissible policies. In this sense, uniform closure captures the syntactic compositionality of the modeling formalism, whereas standard closure captures only its semantic stability.
Unless stated otherwise, all closure results established in this section (e.g., Propositions 3.343.36) refer to the standard notion. The stronger uniform version is discussed separately when explicit constructive transformations $\Phi_{\mathrm{Op}}$ are available, typically at the regular or context-free levels.
Proposition 3.34.
$\mathcal{C}_{0}=\mathcal{C}_{F}=\mathrm{REG}$ is closed under union, intersection, complement, concatenation, and Kleene star.
Proposition 3.35.
$\mathcal{C}_{\mathrm{Stack}}=\mathrm{CFL}$ is closed under union, concatenation, and Kleene star, but not under intersection or complement.
Proposition 3.36.
$\mathcal{C}_{\mathrm{Tape}}=\mathrm{RE}$ is closed under union, concatenation, and Kleene star, but not under complement.
Remark 3.37.
The closure properties stated in Propositions 3.343.36 follow immediately from the classical results of the Chomsky hierarchy. Since these are well-established in the literature, no proofs are included here.
The preceding propositions summarize the standard closure properties enjoyed by the global language families generated under memoryless, finite-memory, stack-augmented and tape-augmented policies. These results follow directly from the classical Chomsky hierarchy and therefore require no additional assumptions about the structure of the underlying controlled automata.
In many applications, however, one is also interested in the stronger question of whether such closure properties can be realized uniformly at the level of the plant structure itself. That is, one asks whether operations such as union or intersection admit canonical automata-level constructions that preserve the control interface and remain compatible with the corresponding policy class.
The next proposition shows that, at the regular level ($\Pi\in\{0,F\}$), this stronger form of uniform closure indeed holds via the standard disjoint-union and synchronous-product constructions.
Proposition 3.38.
For each $\Pi\in\{0,F\}$, the family
$\mathbb{L}_{\Pi}=\{\,L_{\Pi}(S)\mid S\in\mathcal{S}\,\}$
of languages generated by controlled finite automata under memoryless and finite-memory policies is uniformly closed under union and intersection in the sense of Definition 3.32(a).
Proof Sketch.
Fix $\Pi\in\{0,F\}$. Let $S_{1},S_{2}\in\mathcal{S}$ be controlled automata with $L_{i}\in L_{\Pi}(S_{i})$ realized by policies $\pi_{i}\in\Pi$.
(Union). Construct the disjoint-union plant
$S_{\cup}=S_{1}\uplus S_{2},$
whose control and input interfaces coincide with those of $S_{1}$ and $S_{2}$. A policy $\pi_{\cup}$ on $S_{\cup}$ operates by internally simulating $\pi_{1}$ on the $S_{1}$-component and $\pi_{2}$ on the $S_{2}$-component, yielding
$L_{\pi_{\cup}}(S_{\cup})=L_{\pi_{1}}(S_{1})\cup L_{\pi_{2}}(S_{2})=L_{1}\cup L_{2}.$
Since the construction of $S_{\cup}$ depends only on $(S_{1},S_{2})$ and not on the specific policies, the corresponding constructor $\Phi_{\cup}$ realizes uniform closure under union.
(Intersection). For intersection, define the synchronous product
$S_{\cap}=S_{1}\otimes S_{2},$
with composite state space $Q_{1}\times Q_{2}$, shared input alphabet $A$, and joint control space $C_{1}\times C_{2}$. The transition relation is
$((q_{1},q_{2}),a,(c_{1},c_{2}),(q_{1}^{\prime},q_{2}^{\prime}))\in\rightarrow_{\cap}\;\;\text{iff}\;\;(q_{1},a,c_{1},q_{1}^{\prime})\in\rightarrow_{1}\;\text{and}\;(q_{2},a,c_{2},q_{2}^{\prime})\in\rightarrow_{2}.$
The product policy $\pi_{\cap}$ is given by
$\gamma_{\cap}((m_{1},m_{2}),(q_{1},q_{2}))=(\gamma_{1}(m_{1},q_{1}),\,\gamma_{2}(m_{2},q_{2})),$
and
$\eta_{\cap}((m_{1},m_{2}),(q_{1},q_{2}),a,(m_{1}^{\prime},m_{2}^{\prime}))\;\text{iff}\;(m_{i},q_{i},a,m_{i}^{\prime})\in\eta_{i}\text{ for }i=1,2.$
It follows that
$L_{\pi_{\cap}}(S_{\cap})=L_{\pi_{1}}(S_{1})\cap L_{\pi_{2}}(S_{2})=L_{1}\cap L_{2}.$
Thus, intersection is uniformly realizable at the plant level.
In both cases, the constructive mappings $\Phi_{\cup}$ and $\Phi_{\cap}$ are independent of specific policies and depend only on the component plants. Hence, $\mathbb{L}_{\Pi}$ is uniformly closed under union and intersection. ∎
Remark 3.39.
The disjoint-union and synchronous-product constructions on controlled automata preserve the control interface, ensuring that the resulting product plant $S_{1}\otimes S_{2}$ remains controllable by a product policy in the same class $\Pi$.
Proposition 3.40.
Let $\mathcal{S}^{\det}\subseteq\mathcal{S}$ be the class of deterministic controlled finite automata $S=(Q,A,C,\rightarrow,q_{0},F)$ in the sense of Definition 3.21. For each $\Pi\in\{0,F\}$, the family
$\mathbb{L}^{\det}_{\Pi}\;=\;\{\,L_{\Pi}(S)\mid S\in\mathcal{S}^{\det}\,\}$
is uniformly closed under complement in the sense of Definition 3.32(a). More precisely, there is a plant-level constructor
$\Phi_{\complement}:\mathcal{S}^{\det}\to\mathcal{S}^{\det},\qquad\Phi_{\complement}\bigl(Q,A,C,\to,q_{0},F\bigr)\;=\;\bigl(Q,A,C,\to,q_{0},\,Q\setminus F\bigr),$
such that for every policy $\pi\in\Pi$,
$L_{\pi}\bigl(\Phi_{\complement}(S)\bigr)\;=\;A^{*}\setminus L_{\pi}(S).$
Proof Sketch.
Fix a deterministic plant $S=(Q,A,C,\to,q_{0},F)$ and $\Pi\in\{0,F\}$. Determinism means that the relation $\to$ induces a (partial) function $\delta:Q\times A\times C\rightharpoonup Q$; if desired, first complete $S$ by adding a nonaccepting sink so that $\delta$ becomes total—this does not affect languages except for undefined cases, which are mapped to the sink.
Define $S^{\complement}=\Phi_{\complement}(S)$ by keeping the same transition structure and flipping the accepting set to $Q\setminus F$. Consider any policy $\pi\in\Pi$ (memoryless or finite-memory). Because the closed-loop next state is uniquely determined at each step by $(q,a,\gamma(\cdot))$, the unique run of $S$ under $\pi$ on any input word $w\in A^{*}$ visits $F$ at the terminal position iff the unique run of $S^{\complement}$ under the same $\pi$ visits $Q\setminus F$. Thus $w\in L_{\pi}(S)$ iff $w\notin L_{\pi}(S^{\complement})$, yielding $L_{\pi}(S^{\complement})=A^{*}\setminus L_{\pi}(S)$. Since $\Phi_{\complement}$ depends only on the plant (not on $\pi$), this is a uniform plant-level realization of complement on $\mathcal{S}^{\det}$. ∎
Remark 3.41.
Uniform complement at the plant level fails in general for nondeterministic plants: flipping $F$ does not invert each $L_{\pi}(S)$ when multiple runs on the same input/control exist. Determinism (and completion, if needed) ensures a single run per input/control, so, complementing acceptance yields language complement for all $\pi\in\{0,F\}$.
Remark 3.42.
The preceding two propositions show that, at the regular level ($\Pi\in\{0,F\}$), the classical closure properties of regular languages can be lifted to the stronger, model-level notion of uniform closure in Definition 3.32(a). For the stack- and tape- augmented levels ($\Pi\in\{\mathrm{Stack},\mathrm{Tape}\}$), the analogous uniform constructions would have to operate on pushdown- and Turing-like plants and simultaneously account for all admissible control policies on those plants. Developing such plant-level constructors is considerably more involved and lies beyond the scope of this paper. Accordingly, all closure results for $\mathcal{C}_{\mathrm{Stack}}$ and $\mathcal{C}_{\mathrm{Tape}}$ in Propositions 3.353.36 are interpreted in the standard (global) sense.
Emptiness Problem for Controlled Automata on Finite Words
Definition 3.43.
The emptiness problem for a controlled automaton $S$ under a policy class $\Pi$ asks whether there exists a policy $\pi\in\Pi$ such that the language $L_{\pi}(S)$ is nonempty, i.e.,
$\exists\pi\in\Pi\text{ such that }L_{\pi}(S)\neq\emptyset.$
This problem determines whether there exists at least one finite execution accepted by the controlled system under some control strategy.
Proposition 3.44.
The emptiness problem for controlled automata depends on the policy class: it is decidable in polynomial time for $\Pi\in\{0,F\}$, decidable for $\Pi=\textit{Stack}$, and undecidable for $\Pi\in\{\textit{Tape},H\}$.
Proof Sketch.
For $\Pi\in\{0,F\}$, controlled automata reduce to nondeterministic finite automata (NFAs). The emptiness problem for NFAs can be decided in polynomial time via a reachability check over the underlying state graph [Hopcroft2007].
For $\Pi=\textit{Stack}$, the model corresponds to a pushdown automaton, and emptiness is known to be decidable using standard PDA algorithms [Hopcroft2007].
For $\Pi\in\{\textit{Tape},H\}$, the policy mechanism can simulate a Turing machine, and the emptiness problem reduces to the halting problem, which is undecidable [Sipser2013, Hopcroft1979]. ∎
Nondeterministic Languages on Infinite Words
Next, we consider the reactive aspect of a controlled automaton by considering its behavior under an infinite sequence of activity completions as follows.
Definition 3.45.
A controlled Büchi automaton is a 6-tuple $S_{\omega}=(Q,A,C,\rightarrow,Q_{0},F)$ where:
  • $(Q,A,C,\rightarrow,Q_{0})$ is a controlled automaton,
  • $F\subseteq Q$: set of accepting states (visited infinitely often).
Definition 3.46.
Let $S_{\omega}=(Q,A,C,\rightarrow,Q_{0},F)$ be a controlled Büchi automaton and $\pi$ a policy under $S_{\omega}$. The $\omega$-language accepted by $S_{\omega}$ under $\pi$ is defined as:
[…]
where $\text{Inf}(\rho)$ denotes the set of states that occur infinitely often in the projection of $\rho$ onto $Q$.
Proposition 3.47.
Define the following $\omega$-language classes:
$L^{\omega}_{0}(S_{\omega})$ $=\{L^{\omega}_{\pi}(S_{\omega})\mid\pi\text{ is a memoryless policy}\},$
$L^{\omega}_{F}(S_{\omega})$ $=\{L^{\omega}_{\pi}(S_{\omega})\mid\pi\text{ is a finite-memory policy}\},$
$L^{\omega}_{\text{Stack}}(S_{\omega})$ $=\{L^{\omega}_{\pi}(S_{\omega})\mid\pi\text{ is a stack-augmented policy}\},$
$L^{\omega}_{\text{Tape}}(S_{\omega})$ $=\{L^{\omega}_{\pi}(S_{\omega})\mid\pi\text{ is a tape-augmented policy}\},$
$L^{\omega}_{H}(S_{\omega})$ $=\{L^{\omega}_{\pi}(S_{\omega})\mid\pi\text{ is a history-dependent policy}\}.$
Then, we have:
$L^{\omega}_{0}(S_{\omega})\subset L^{\omega}_{F}(S_{\omega})\subset L^{\omega}_{\text{Stack}}(S_{\omega})\subset L^{\omega}_{\text{Tape}}(S_{\omega})\subset L^{\omega}_{H}(S_{\omega})$
Definition 3.48.
Let $S_{\omega}=(Q,A,C,\rightarrow,Q_{0},F)$ and $S_{\omega}^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},\rightarrow^{\prime},Q^{\prime}_{0},F^{\prime})$ be two equivalent controlled Büchi automata with the same activity alphabet and set of control actions (i.e., $A=A^{\prime}$ and $C=C^{\prime}$). Let $(Q,A,C,\rightarrow,Q_{0})$ and $(Q^{\prime},A^{\prime},C^{\prime},\rightarrow^{\prime},Q^{\prime}_{0})$ be two equivalent controlled automata under a bisimulation $\gamma$ as in Definition  3.5. Then, $S_{\omega}$ and $S^{\prime}_{\omega}$ are said to be equivalent if:
  • for any $q\in F$, there exists $q^{\prime}\in F^{\prime}$ where $(q,q^{\prime})\in\gamma$; also for any $q^{\prime}\in F^{\prime}$, there exists $q\in F$ where $(q^{\prime\prime},q)\in\gamma$.
Proposition 3.49.
Let $S_{\omega}$ and $S^{\prime}_{\omega}$ be two equivalent controlled Büchi automata with accepting states as in Definition 3.48. Consider the classes of languages as defined in Proposition 3.47. Then, we have:
$L^{\omega}_{0}(S_{\omega})$ $=L^{\omega}_{0}(S^{\prime}_{\omega})$
$L^{\omega}_{F}(S_{\omega})$ $=L^{\omega}_{F}(S^{\prime}_{\omega})$
$L^{\omega}_{\text{Stack}}(S_{\omega})$ $=L^{\omega}_{\text{Stack}}(S^{\prime}_{\omega})$
$L^{\omega}_{\text{Tape}}(S_{\omega})$ $=L^{\omega}_{\text{Tape}}(S^{\prime}_{\omega})$
$L^{\omega}_{H}(S_{\omega})$ $=L^{\omega}_{H}(S^{\prime}_{\omega})$
Also, for equivalent polices $\pi$ and $\pi^{\prime}$ of $S_{\omega}$ and $S^{\prime}_{\omega}$, respectively, as in Definition 3.16, we have:
$L^{\omega}_{\pi,0}(S_{\omega})$ $=L^{\omega}_{\pi^{\prime},0}(S^{\prime}_{\omega})$
$L^{\omega}_{\pi,F}(S_{\omega})$ $=L^{\omega}_{\pi^{\prime},F}(S^{\prime}_{\omega})$
$L^{\omega}_{\pi,\text{Stack}}(S_{\omega})$ $=L^{\omega}_{\pi^{\prime},\text{Stack}}(S^{\prime}_{\omega})$
$L^{\omega}_{\pi,\text{Tape}}(S_{\omega})$ $=L^{\omega}_{\pi^{\prime},\text{Tape}}(S^{\prime}_{\omega})$
$L^{\omega}_{\pi,H}(S_{\omega})$ $=L^{\omega}_{\pi^{\prime},H}(S^{\prime}_{\omega})$
Global Language Families for Controlled Büchi Automata
We now summarize the structure and properties of global language families defined by controlled Büchi automata, extending the hierarchy developed for controlled automata on finite words.
Definition 3.50.
Let $\mathcal{S}_{\omega}$ be the set of all controlled Büchi automata. For each class of control policies $\Pi\in\{0,F,\text{Stack},\text{Tape},H\}$, define:
$\mathbb{L}^{\omega}_{\Pi}:=\left\{L^{\omega}_{\pi}(S_{\omega})\mid\pi\in\Pi,\ S_{\omega}\in\mathcal{S}_{\omega}\right\}\subseteq\mathcal{P}(2^{A^{\omega}})$
where each $L^{\omega}_{\pi}(S_{\omega})$ is the set of infinite words $w\in A^{\omega}$ generated by $S_{\omega}$ under policy $\pi$ that satisfy the Büchi acceptance condition.
Interpretation
Each $\mathbb{L}^{\omega}_{\Pi}$ is a set of infinite-word language classes, indexed by the automaton $S_{\omega}$ and parameterized by the control policy class $\Pi$. Just as in the finite-word setting, this forms a second-order structure:
$\mathbb{L}^{\omega}_{\Pi}\subseteq\mathcal{P}(2^{A^{\omega}})$
where each element is a set of infinite-word languages.
Hierarchy
As in the finite case, there is a strict hierarchy of global language families:
Theorem 3.51.
For the $\omega$-language classes defined in Definition 3.50, we have
$\mathbb{L}^{\omega}_{0}\subset\mathbb{L}^{\omega}_{F}\subset\mathbb{L}^{\omega}_{\text{Stack}}\subset\mathbb{L}^{\omega}_{\text{Tape}}=\mathbb{L}^{\omega}_{H^{\text{comp}}}\subseteq\mathbb{L}^{\omega}_{H}\,.$
Proof Sketch..
The argument parallels that of Theorem 3.29 for finite-word languages, adapted to $\omega$-languages under Büchi acceptance. Each inclusion corresponds to a strictly increasing level of controller power: memoryless $\subset$ finite-memory $\subset$ stack-augmented $\subset$ tape-augmented == computable history-dependent $\subset$ arbitrary history-dependent. Strictness follows from analogous examples where finite memory distinguishes finite phases, stacks enforce well-nested obligations, and tapes simulate arbitrary computations. Hence,
$\mathbb{L}^{\omega}_{0}\subset\mathbb{L}^{\omega}_{F}\subset\mathbb{L}^{\omega}_{\text{Stack}}\subset\mathbb{L}^{\omega}_{\text{Tape}}=\mathbb{L}^{\omega}_{H^{\text{comp}}}\subseteq\mathbb{L}^{\omega}_{H}\,.$
Lemma 3.52.
There exists a controlled Büchi automaton $S_{\omega}$ such that $\mathbb{L}^{\omega}_{F}(S_{\omega})$ contains a countably infinite family of distinct $\omega$-regular languages.
Proof Sketch..
This construction is the $\omega$-word analogue of Lemma 3.30. We build a controlled Büchi automaton $S_{\omega}$ over $\Sigma=\{a,b\}$ whose transitions allow two phases: an $a$-phase of length $k$, followed by a $b$-reset, repeating forever. A finite-memory controller can store a counter modulo $k$ and permit exactly $k$ $a$-steps before forcing a $b$, then repeat. Under such a controller, the accepted $\omega$-language is $(a^{k}b)^{\omega}$, which is $\omega$-regular.
By varying $k\geq 1$, we obtain a countably infinite family $\{(a^{k}b)^{\omega}\mid k\geq 1\}$ of pairwise distinct $\omega$-regular languages, each realizable by some finite-memory policy on the same plant $S_{\omega}$. Thus $L_{F}^{\omega}(S_{\omega})$ contains a countably infinite family of distinct $\omega$-regular languages. ∎
Proposition 3.53.
There exists a countably infinite family of pairwise distinct omega\ omega-regular languages that is not equal to $L^{\omega}_{F}(S_{\omega})$ for any single controlled Büchi automaton $S_{\omega}$.
Proof Sketch..
This argument is the $\omega$-word analogue of Proposition 3.31. For each integer $k\geq 1$, define the $\omega$-regular language $W_{k}=(a^{k}b)^{\omega}$, i.e., the ultimately periodic infinite word consisting of blocks of $a^{k}$ followed by $b$, repeated forever. For distinct $k$, these $W_{k}$ are distinct $\omega$-regular languages.
Let
$\mathcal{F}^{\omega}\;=\;\{\,W_{p}\mid p\text{ prime}\,\}\;=\;\{\,(a^{p}b)^{\omega}\mid p\text{ prime}\,\}.$
Clearly $\mathcal{F}^{\omega}$ is a countably infinite family.
Suppose, toward contradiction, that there exists a controlled Büchi automaton $S_{\omega}$ such that $L_{F}^{\omega}(S_{\omega})=\mathcal{F}^{\omega}$, i.e., that by varying finite-memory policies on $S_{\omega}$, we obtain exactly the languages in $\mathcal{F}^{\omega}$ and no others.
Fix a prime $p$. A finite-memory controller that enforces $W_{p}=(a^{p}b)^{\omega}$ can be seen as a finite-state machine whose cycle has period $p$: it allows exactly $p$ $a$’s, then a $b$, then repeats. But from such a controller, we can build another finite-memory controller that simply iterates that internal cycle $p$ times before emitting $b$. The resulting policy produces $(a^{p^{2}}b)^{\omega}=W_{p^{2}}$. Since $p^{2}$ is not prime, $W_{p^{2}}\notin\mathcal{F}^{\omega}$ by definition. Yet $W_{p^{2}}$ would still lie in $L_{F}^{\omega}(S_{\omega})$ because the modified controller is still finite-memory.
This contradicts the assumption that $L_{F}^{\omega}(S_{\omega})$ equals $\mathcal{F}^{\omega}$. Hence, there is a countably infinite family of distinct $\omega$-regular languages (namely $\mathcal{F}^{\omega}$) that is not equal to $L_{F}^{\omega}(S_{\omega})$ for any $S_{\omega}$. ∎
Remark 3.54.
These results parallel Lemma 3.30 and Proposition 3.31 for finite words. The class $\mathbb{L}^{\omega}_{F}$ is strictly between $\mathbb{L}^{\omega}_{0}$ and $\mathbb{L}^{\omega}_{\mathrm{Stack}}$ and inherits the expressive power and closure properties of $\omega$-regular languages.
The preceding results establish the structural hierarchy of the $\omega$-language families and illustrate how finite-memory controllers give rise to countably infinite yet well-characterized $\omega$-regular languages. We now turn to further properties of these language classes, focusing on their algebraic behavior under standard operations and their closure characteristics across policy types. The following four propositions formally capture these closure relations.
The closure properties established in the preceding subsection for finite languages naturally extend to the $\omega$-language setting once the acceptance semantics of controlled Büchi automata are taken into account. In particular, the same distinction between uniform (constructive) and standard (semantic) closure remains meaningful, as the basic operations on $\omega$-languages—such as union, intersection, concatenation with finite prefixes, and Büchi projection—can likewise be viewed either as syntactic transformations on plant structures or as semantic operations on the induced language families. To make this correspondence precise, we now introduce the formal definition of uniform and standard closure under $\omega$-language operations.
Closure Properties of Controlled $\omega$-Language Families
The closure analysis naturally extends to the $\omega$-language setting, where infinite behaviors are captured by controlled Büchi automata. We formalize the corresponding closure notions for these models and relate them to the classical $\omega$-language hierarchy established by Büchi, McNaughton, and Landweber.
Definition 3.55 (Uniform and Standard Closure under $\omega$-Language Operations).
Let $\mathcal{S}_{\omega}$ denote the class of controlled Büchi automata (plants accepting $\omega$-languages), and for each policy class $\Pi\in\{0,F,\mathrm{Stack},\mathrm{Tape}\}$ let $L^{\omega}_{\Pi}(S)$ denote the $\omega$-languages realizable by policies $\pi\in\Pi$ on a plant $S\in\mathcal{S}_{\omega}$.
(a) Uniform Closure. The family
$\mathbb{L}^{\omega}_{\Pi}=\{\,L^{\omega}_{\Pi}(S)\mid S\in\mathcal{S}_{\omega}\,\}$
is said to be uniformly closed under a $k$-ary $\omega$-language operation $Op:(P(A^{\omega}))^{k}\!\to\!P(A^{\omega})$ if there exists a computable constructor
$\Phi^{\omega}_{Op}:\mathcal{S}_{\omega}^{k}\longrightarrow\mathcal{S}_{\omega}$
such that for all plants $S_{1},\dots,S_{k}\in\mathcal{S}_{\omega}$ and all $L_{i}\in L^{\omega}_{\Pi}(S_{i})$ ($1\!\leq\!i\!\leq\!k$),
$Op(L_{1},\dots,L_{k})\in L^{\omega}_{\Pi}\bigl(\Phi^{\omega}_{Op}(S_{1},\dots,S_{k})\bigr).$
Thus, the same syntactic transformation $\Phi^{\omega}_{Op}$ must realize the operation $Op$ for all admissible $\omega$-languages in the class.
(b) Standard (Pointwise) Closure. The global $\omega$-language family
$\mathcal{C}^{\omega}_{\Pi}=\bigcup_{S\in\mathcal{S}_{\omega}}L^{\omega}_{\Pi}(S)$
is closed under $Op$ if for all $L_{1},\dots,L_{k}\in\mathcal{C}^{\omega}_{\Pi}$,
$Op(L_{1},\dots,L_{k})\in\mathcal{C}^{\omega}_{\Pi}.$
Equivalently, for every collection of plants $S_{1},\dots,S_{k}\in\mathcal{S}_{\omega}$ and policies $\pi_{1},\dots,\pi_{k}\in\Pi$, there exist a plant $S^{\prime}\in\mathcal{S}_{\omega}$ and a policy $\pi^{\prime}\in\Pi$ such that
$Op\!\bigl(L^{\omega}_{\pi_{1}}(S_{1}),\dots,L^{\omega}_{\pi_{k}}(S_{k})\bigr)=L^{\omega}_{\pi^{\prime}}(S^{\prime}).$
In contrast with the uniform case, no explicit structural correspondence between $S^{\prime}$ and the original plants is required.
Having introduced the notion of closure for controlled Büchi automata, we now summarize the corresponding algebraic properties of the induced $\omega$-language families. As in the finite-word setting, these results characterize how the expressive classes $\mathcal{C}^{\omega}_{0}$, $\mathcal{C}^{\omega}_{F}$, $\mathcal{C}^{\omega}_{\mathrm{Stack}}$, and $\mathcal{C}^{\omega}_{\mathrm{Tape}}$ behave under fundamental operations such as union, intersection, and complement. Two complementary notions of closure are relevant here: the standard (language-level) closure, which concerns the semantic stability of these families under the operations, and the uniform (model-level) closure, which requires the existence of explicit constructive transformations on controlled Büchi automata that realize those operations uniformly across all admissible policies. Unless stated otherwise, the propositions below refer to the standard notion, while the uniform counterpart is discussed later when constructive realizations can be defined explicitly.
Remark 3.56.
The distinction between standard and uniform closure introduced above applies equally to the $\omega$-language setting. In this context, uniform closure corresponds to the existence of explicit constructive transformations $\Phi_{\mathrm{Op}}^{\omega}$ on controlled Büchi automata that realize each operation at the model level. Unless stated otherwise, the closure results in this subsection are stated with respect to the standard notion.
Proposition 3.57.
$\mathcal{C}^{\omega}_{0}=\mathcal{C}^{\omega}_{F}=\mathrm{REG}^{\omega}$ is closed under finite union, intersection, and complementation.
Proof Sketch.
By Büchi’s theorem [Buchi1962] and McNaughton’s determinization result [McNaughton1966], every $\omega$-regular language is accepted by a deterministic Muller automaton, and deterministic Muller automata are effectively closed under union, intersection, and complementation. Hence, $\mathrm{REG}^{\omega}$ enjoys full Boolean closure. ∎
Proposition 3.58.
$\mathcal{C}^{\omega}_{\mathrm{Stack}}=\mathrm{CFL}^{\omega}$ is closed under finite union and under intersection with $\omega$-regular languages, but not under general intersection or complementation.
Proof Sketch.
Closure under union follows from the characterization of $\omega$-CFLs as finite unions of languages of the form $UV^{\omega}$ with $U,V$ context-free, as shown by Cohen and Gold [CohenGold1977]. Non-closure under intersection and complement is classical: there exist $\omega$-CFLs $L_{1},L_{2}$ such that $L_{1}\cap L_{2}$ is not $\omega$-context-free, implying, by De Morgan’s laws, failure of closure under complement (see also Staiger [Staiger1997] and Finkel [Finkel1997]). ∎
Proposition 3.59.
$\mathcal{C}^{\omega}_{\mathrm{Tape}}=\mathrm{RE}^{\omega}$ is closed under finite union and intersection, but not under complementation.
Proof Sketch.
The class $\mathrm{RE}^{\omega}$ coincides with the family of analytic ($\Sigma^{1}_{1}$) subsets of $A^{\omega}$, known to be closed under finite union and intersection but not under complement (Kechris [Kechris1995], Moschovakis [Moschovakis2009]). ∎
The preceding proposition summarizes the standard closure properties of the global $\omega$-language families generated by controlled Büchi automata. These results rely solely on the expressive power of the underlying acceptance model and do not require any structural relationship between the plants that realize the participating languages.
For compositional reasoning, however, one is often interested in a stronger notion: whether the same operations can be implemented uniformly by plant-level constructions that work for all memoryless and finite-memory policies. In the finite-word setting, this was achieved using disjoint union and the synchronous product of plants. The following proposition shows that these constructions extend naturally to controlled Büchi automata, yielding uniform closure under finite union and intersection at the $\omega$-regular level.
Proposition 3.60.
For each $\Pi\in\{0,F\}$, the family
$\mathbb{L}^{\omega}_{\Pi}=\{\,L^{\omega}_{\Pi}(S)\mid S\in\mathcal{S}_{\omega}\,\}$
of $\omega$-languages generated by controlled Büchi automata under memoryless and finite-memory policies is uniformly closed under finite union and intersection in the sense of Definition 3.55(a).
Proof Sketch.
Fix $\Pi\in\{0,F\}$ and let $S_{1},S_{2}\in\mathcal{S}_{\omega}$ be controlled Büchi automata with $L_{i}=L^{\omega}_{\pi_{i}}(S_{i})$ realized by policies $\pi_{i}\in\Pi$.
(Union). Construct the disjoint-union plant
$S_{\cup}=S_{1}\uplus S_{2},$
whose input alphabet and control interface coincide with those of $S_{1}$ and $S_{2}$. A policy $\pi_{\cup}$ on $S_{\cup}$ operates by internally simulating $\pi_{1}$ on the $S_{1}$-component and $\pi_{2}$ on the $S_{2}$-component, with the acceptance set $F_{\cup}=F_{1}\cup F_{2}$. Hence,
$L^{\omega}_{\pi_{\cup}}(S_{\cup})=L^{\omega}_{\pi_{1}}(S_{1})\cup L^{\omega}_{\pi_{2}}(S_{2})=L_{1}\cup L_{2}.$
Since the construction of $S_{\cup}$ depends only on the plants $(S_{1},S_{2})$, this defines a uniform constructor $\Phi^{\omega}_{\cup}$ for union.
(Intersection). For intersection, define the synchronous product
$S_{\cap}=S_{1}\otimes S_{2}$
with state space $Q_{1}\times Q_{2}$, shared input alphabet $A$, and joint control space $C_{1}\times C_{2}$. The transition relation is given by
$((q_{1},q_{2}),a,(c_{1},c_{2}),(q_{1}^{\prime},q_{2}^{\prime}))\in\rightarrow_{\cap}\quad\text{iff}\quad(q_{1},a,c_{1},q_{1}^{\prime})\in\rightarrow_{1}\text{ and }(q_{2},a,c_{2},q_{2}^{\prime})\in\rightarrow_{2}.$
The product policy $\pi_{\cap}$ acts componentwise,
$\gamma_{\cap}((m_{1},m_{2}),(q_{1},q_{2}))=(\gamma_{1}(m_{1},q_{1}),\,\gamma_{2}(m_{2},q_{2})),$
and the Büchi acceptance set is
$F_{\cap}=F_{1}\times F_{2}.$
Then,
$L^{\omega}_{\pi_{\cap}}(S_{\cap})=L^{\omega}_{\pi_{1}}(S_{1})\cap L^{\omega}_{\pi_{2}}(S_{2})=L_{1}\cap L_{2}.$
The plant $S_{\cap}$ depends only on $S_{1}$ and $S_{2}$, and the policy $\pi_{\cap}$ remains within the same class $\Pi$, demonstrating uniform closure under intersection.
In both cases, the constructors $\Phi^{\omega}_{\cup}$ and $\Phi^{\omega}_{\cap}$ are independent of the specific policies and depend only on the plant structures. Hence, $\mathbb{L}^{\omega}_{\Pi}$ is uniformly closed under union and intersection at the $\omega$-regular level. ∎
Remark 3.61.
Disjoint-union and synchronous-product constructions on controlled Büchi automata preserve the control interface, ensuring that the resulting product plant $S_{1}\otimes S_{2}$ remains controllable by a product policy in the same class $\Pi$.
Proposition 3.62.
Let $\mathcal{S}_{\omega}^{\mathrm{det}}$ be the class of deterministic controlled $\omega$-automata $S=(Q,A,C,\to,q_{0},\mathsf{Acc})$ where $\mathsf{Acc}$ is one of: parity, Muller, Rabin, or Streett. For each $\Pi\in\{0,F\}$, the family
$\mathbb{L}^{\omega,\mathrm{det}}_{\Pi}=\bigl\{\,L^{\omega}_{\Pi}(S)\;\bigm|\;S\in\mathcal{S}_{\omega}^{\mathrm{det}}\,\bigr\}$
is uniformly closed under complement. More precisely, there is a plant-level constructor
$\Phi^{\omega}_{\complement}:\mathcal{S}_{\omega}^{\mathrm{det}}\longrightarrow\mathcal{S}_{\omega}^{\mathrm{det}}$
that keeps $(Q,A,C,\to,q_{0})$ unchanged and replaces $\mathsf{Acc}$ by its dual:
  • Parity: priorities $\Omega:Q\to\{0,\dots,d\}$ become $\Omega^{\prime}(q)=\Omega(q)+1$ (i.e., flip even/odd).
  • Muller: $\mathcal{F}\subseteq 2^{Q}$ becomes $\mathcal{F}^{\prime}=2^{Q}\setminus\mathcal{F}$.
  • Rabin/Streett: swap each pair $(E_{i},F_{i})$ to $(F_{i},E_{i})$, turning Rabin $\leftrightarrow$ Streett.
Then for every policy $\pi\in\Pi$,
$L^{\omega}_{\pi}\!\bigl(\Phi^{\omega}_{\complement}(S)\bigr)=A^{\omega}\setminus L^{\omega}_{\pi}(S).$
Proof Sketch.
Fix $S\in\mathcal{S}_{\omega}^{\mathrm{det}}$ and $\pi\in\Pi$. Determinism ensures that for any input word there is a unique run, hence a unique set $\mathrm{Inf}(\rho)\subseteq Q$ of states visited infinitely often under $\pi$. Each listed dualization (parity flip, Muller complement, Rabin/Streett swap) is the classical acceptance dual that accepts exactly when the original rejects on the same run. Since $\Phi^{\omega}_{\complement}$ depends only on $S$ (not on $\pi$), this is a uniform plant-level construction yielding language complement for all $\pi\in\{0,F\}$. ∎
Remark 3.63.
We deliberately exclude deterministic Büchi acceptance: deterministic Büchi automata are not closed under complement. If one models plants with deterministic parity/Muller/Rabin/Streett, the above uniform complement applies without changing the transition graph.
Remark 3.64.
If, in addition, $\mathcal{S}_{\omega}$ is restricted to controlled deterministic parity automata (or any equivalent complement-closed $\omega$-regular formalism), then the same framework yields uniform closure under complementation as well, by replacing the parity condition with its dual on a fixed deterministic plant. Since we do not impose this structural restriction in general, Proposition 3.60 is stated only for union and intersection. It is well known that deterministic Büchi automata are not closed under complement, whereas deterministic parity or Rabin automata are. Hence, uniform complementation requires augmenting the acceptance structure beyond the Büchi property (e.g., to parity, Rabin/Street, or Muller conditions [Safra1988, Rabin1969, FrancezPnueli1980, Muller1963]).
Remark 3.65.
The results show that the standard closure properties of $\omega$-regular languages (as established by Büchi [Buchi1962], McNaughton [McNaughton1966], and Landweber [Landweber1969]) remain valid at the language level for controlled Büchi automata. However, uniform syntactic realizations exist only in the regular (deterministic) fragment. This highlights the same hierarchy observed in the finite setting: uniform (model-level) closure becomes strictly weaker than standard (semantic) closure once non-determinism or infinite-memory control is introduced.
Remark 3.66.
The preceding proposition shows that, at the $\omega$-regular level ($\Pi\in\{0,F\}$), the classical closure constructions for $\omega$-regular languages can be lifted to the stronger, model-level notion of uniform closure in Definition 3.55(a). For the stack- and tape-augmented levels ($\Pi\in\{\mathrm{Stack},\mathrm{Tape}\}$), extending these uniform constructions to pushdown- and Turing-like Büchi plants would require substantially more involved plant-level transformations, and is not pursued here. Accordingly, the closure properties for $\mathcal{C}^{\omega}_{\mathrm{Stack}}$ and $\mathcal{C}^{\omega}_{\mathrm{Tape}}$ stated in Propositions 3.583.59 are interpreted in the standard (global) sense.
Emptiness Problem for Controlled Büchi Automata
Definition 3.67.
This problem determines whether there exists at least one infinite execution accepted by the controlled Büchi automaton under some control strategy. The emptiness problem for controlled Büchi automata $S_{\omega}\in\mathcal{S}_{\omega}$ asks whether there exists a policy $\pi\in\Pi$ such that the infinite-word language $L_{\pi}^{\omega}(S_{\omega})$ is nonempty:
$\exists\pi\in\Pi\text{ such that }L^{\omega}_{\pi}(S_{\omega})\neq\emptyset.$
Proposition 3.68.
For the emptiness problem for controlled Büchi automata, we have:
  • For $\Pi\in\{0,F\}$, the problem is PSPACE-complete and decidable.
  • For $\Pi=\textit{Stack}$, the problem remains decidable using pushdown Büchi automata techniques.
  • For $\Pi\in\{\textit{Tape},H\}$, the problem is undecidable.
Proof Sketch.
For $\Pi\in\{0,F\}$, controlled Büchi automata reduce to classical nondeterministic Büchi automata. The emptiness problem for these automata is PSPACE-complete. It can be decided using graph-search algorithms that detect reachable accepting cycles, such as nested depth-first search (DFS) [baier2008principles, Vardi1985, Buchi1962].
For $\Pi=\textit{Stack}$, the model corresponds to a pushdown Büchi automaton. The emptiness problem remains decidable via saturation techniques and automata-theoretic constructions developed for pushdown systems [Esparza1997PDABuchi, Walukiewicz2001, Bouajjani2000]. Decidability assumes a finite stack alphabet and a finite-state plant; the Büchi set is evaluated on the standard product pushdown system.
For $\Pi\in\{\textit{Tape},H\}$, the control mechanism can simulate a Turing machine on infinite inputs, and the emptiness problem becomes equivalent to the nonemptiness problem for $\omega$-languages accepted by Turing machines, which is undecidable [Sipser2013, Hopcroft1979]. ∎
Remark 3.69.
The PSPACE-completeness for $\Pi\in\{0,F\}$ follows directly from the classical result on the nonemptiness problem for nondeterministic Büchi automata. In this case, the complexity arises from the need to explore accepting strongly connected components. Still, it remains tractable within PSPACE due to the linear-space nature of DFS-based algorithms.
4 Probabilistic Models
The following models are extensions of controlled activity networks where nondeterminacy is specified probabilistically. This is accomplished by assigning probabilities to various instantaneous activities.
Model Structure
Definition 4.1.
A controlled probabilistic activity network is a tuple $(K,IP)$ where:
  • $K=(P,IA,TA,C,IG,OG,IR,IOR,TOR)$ is a controlled activity network,
  • $IP:{\mathcal{N}}^{n}\times IA\longrightarrow[0,\;\;1]$ is the instantaneous activity probability distribution, where $n=|P|$ and $IP$ is a probability distribution over ${\mathcal{N}}^{n}\times IA$.
Model Behavior
The behavior of the above model is similar to that of a controlled activity network, except that when there is more than one enabled instantaneous activity in an unstable marking, the choice of which activity completes first is made probabilistically. More specifically, let $L$ be a controlled probabilistic activity network as in Definition  4.1 Suppose, $L$ is in an unstable marking $\mu$. Let $A^{\prime}$ be the set of enabled instantaneous activities of $L$ in $\mu$. Then, $a\in A^{\prime}$ completes with probability $\alpha$, where
$\alpha=\frac{IP(\mu,a)}{\sum_{a^{\prime}\in A^{\prime}}IP(\mu,a^{\prime})}.$
The above summarizes the behavior of a controlled probabilistic activity network.
Semantic Models
To study this behavior more formally, we need to define the notion of a controlled probabilistic automaton.
Definition 4.2.
A controlled probabilistic automaton is a 5-tuple $(Q,A,C,P,Q_{0})$ where:
  • $Q$ is a set of states,
  • AA is the activity alphabet,
  • $C$ is the control action alphabet,
  • $P:Q\times A\times C\times Q\to[0,1]$ is the probabilistic transition function, such that:
    $\sum_{q^{\prime}\in Q}P(q,a,c,q^{\prime})=1\quad\text{for all }(q,a,c)\in Q\times A\times C,$
  • $Q_{0}$ is the initial state distribution which is a probability distribution over $Q$.
For $a\in A$ and $q^{\prime}\in Q, qq^{\prime}$ is said to be immediately reachable from $q$ under $a$ with probability $\alpha$, if $P(q,a,q^{\prime})=\alpha$.
Bismulation
We now present a notion of equivalence for controlled probabilistic automata which is similar to the one proposed for probabilistic automata [Larsen1991].
Definition 4.3.
Let $U=(Q,A,C,P,Q_{0})$ and $U^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},P^{\prime},Q^{\prime}_{0})$ be two probabilistic automata with the same activity alphabet and set of control actions(i.e., $A=A^{\prime}$ and $C=C^{\prime}$). $U$ and $U^{\prime}$ are said to be equivalent if there exists a symmetric binary relation $\gamma$ on $Q\cup Q^{\prime}$ such that:
  • for any $q\in Q$, there exists a $q^{\prime}\in Q^{\prime}$ such that $(q,q^{\prime})\in\gamma$; also, for any $q^{\prime}\in Q^{\prime}$, there exists a $q\in Q$ such that $(q^{\prime},q)\in\gamma$,
  • for any $q_{0}\in Q$ and $q^{\prime}_{0}\in Q^{\prime}$ such that $(q_{0},q^{\prime}_{0})\in\gamma$,
    $\sum_{(q,q^{\prime}_{0})\in\gamma}Q_{0}(q)=\sum_{(q^{\prime},q_{0})\in\gamma}Q^{\prime}_{0}(q^{\prime}),$
  • for any $a\in A$, $c\in C$, $q_{1},q_{2}\in Q$, and $q^{\prime}_{1},q^{\prime}_{2}\in Q^{\prime}$ such that $(q_{1},q^{\prime}_{1})\in\gamma$ and $(q_{2},q^{\prime}_{2})\in\gamma$,
    $\sum_{(q,q^{\prime}_{2})\in\gamma}P(q_{1},a,c,q)=\sum_{(q^{\prime},q_{2})\in\gamma}P^{\prime}(q^{\prime}_{1},a,c,q^{\prime}).$
$\gamma$ above is said to be a bisimulation between $U$ and $^{\prime}. UU$ and $U^{\prime}$ are isomorphic if $\gamma$ is a bijection.
Proposition 4.4.
Let $\mathcal{E_{U}}$ denote a relation on the set of all controlled probabilistic automata such that $(U_{1},U_{2})\in\mathcal{E_{U}}$ if and only if $U_{1}$ and $U_{2}$ are equivalent controlled probabilistic automata in the sense of Definition  4.3. Then $\mathcal{E_{U}}$ will be an equivalence relation.
Operational Semantics
The behavior of a probabilistic activity network may now be formalized as follows.
Definition 4.5.
Let $(L,\mu_{0})$ denote a controlled probabilistic activity network $L$ with an initial marking $\mu_{0}$ where $L$ is defined as in Definition  4.1. $(L,\mu_{0})$ is said to realize a probabilistic automaton $U=(Q,A,C,P,Q_{0})$ where:
  • $Q$ is the set of all stable markings of $L$ which are reachable from $\mu_{0}$ and a state $\Delta$ if, in $L$, an infinite sequence of instantaneous activities can be completed in a marking reachable from $\mu_{0}$,
  • AA is the set of timed activities of $L$,
  • $C$ is the set of control actions of $L$,
  • For any $\mu,\mu^{\prime}\in Q$, $a\in A$ and $c\in C$ such that $a$ is enabled in $\mu$, $P(\mu,a,c,\mu^{\prime})$ is the probability that, in $L$, $\mu^{\prime}$ is the next stable marking to be reached upon completion of $a$ in $\mu$ via $c$; $P(\mu,a,c,\Delta)$ is the probability that, in $L$, a sequence of activities $ax$ completes in $\mu$, where $x$ is an infinite sequence of instantaneous activities that can be completed after the completion of timed activity $a$ via $c$,
  • For any $\mu\in Q$, $Q_{0}(\mu)$ is the probability that, in $L$, $\mu$ is reached upon completion of a (possibly an empty) string of instantaneous activities in $\mu_{0}$; $Q_{0}(\Delta)$ is the probability that, in $L$, an infinite sequence of instantaneous activities completes in $\mu_{0}$.
A notion of equivalence for controlled probabilistic activity networks may now be given as follows.
Definition 4.6.
Two controlled probabilistic activity networks are equivalent if they realize equivalent probabilistic automata.
Proposition 4.7.
Let $\mathcal{E_{L}}$ denote a relation on the set of all controlled probabilistic activity networks such that $(L_{1},L_{2})\in\mathcal{E_{L}}$ if and only if $L_{1}$ and $L_{2}$ are equivalent controlled probabilistic activity networks in the sense of Definition  4.6. Then $\mathcal{E_{L}}$ will be an equivalence relation.
Policy Types
We now consider the notion of policy in a controlled probabilistic automaton. A policy may be deterministic or probabilistic, depending on whether control actions are defined deterministically or probabilistically. In the previous section, we defined four types of policies, which were deterministic. We can adopt similar policy types in this section as well. However, we introduce some additional probabilistic policy types as follows.
Definition 4.8.
Let $(Q,A,C,P,Q_{0})$ be a controlled probabilistic automaton. A (history-based) probabilistic policy $\pi$ is defined as:
$\pi:(Q\times A)^{*}\to Dist(C),$
where $Dist(C)$ denotes the set of all probability distributions on the set of control actions $C$. $\pi$ is called a memoryless probabilstic policy if $\pi:Q\to Dist(C)$.
Definition 4.9.
Let $U=(Q,A,C,P,Q_{0})$ and $M^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},P^{\prime},Q^{\prime}_{0})$ be two equivalent controlled automata with the same activity alphabet and set of control actions (i.e., $A=A^{\prime}$ and $C=C^{\prime}$). Let $\gamma$ be the corresponding bisimulation relation. Two (history-dependent) policies $\pi$ and $\pi^{\prime}$ in $U$ and $U^{\prime}$, respectively, are said to be equivalent under $\gamma$ if:
  • For runs
    $\rho=(q_{0},a_{0})(q_{1},a_{1})\cdots(q_{n-1},a_{n-1})q_{n},\;\rho^{\prime}=(q^{\prime}_{0},a_{0})(q^{\prime}_{1},a_{1})\cdots(q^{\prime}_{n-1},a_{n-1})q^{\prime}_{n},$
    in $U$ and $U^{\prime}$, respectively, such that $(q_{i},q^{\prime}_{i})\in\gamma$ for $0\leq i\leq n$, the following holds:
    $\pi^{\prime}(h^{\prime}_{i})=\pi(h_{i}),\quad\text{where}\quad\begin{aligned} h_{i}&=(q_{0},a_{0})\ldots(q_{i-1},a_{i-1}),\\ h^{\prime}_{i}&=(q^{\prime}_{0},a_{0})\ldots(q^{\prime}_{i-1},a_{i-1}),\end{aligned}\quad 0<i\leq n.$
Definition 4.10.
A finite-memory probabilistic policy is a tuple $\pi=(M,m_{0},\delta,\gamma)$ where:
  • $M$: finite set of memory states,
  • $m_{0}\in M$: initial memory state,
  • $\delta:M\times(Q\times A)\to M$: memory update function,
  • $\gamma:M\to Dist(C)$: probabilistic control function.
Definition 4.11.
A stack-augmented probabilistic policy is a tuple $\pi=(\Gamma,\bot,m_{0},\delta,\gamma)$ where:
  • $\Gamma$: finite stack alphabet with bottom symbol $\bot$,
  • $m_{0}$: initial configuration (state and stack),
  • $\delta:\Gamma\times Q\times A\rightarrow\Gamma^{*}$: stack update function depending only on the top stack symbol,
  • $\gamma:\Gamma\to Dist(C)$: probabilistic control output function depending only on the top stack symbol.
Definition 4.12.
A tape-augmented probabilistic policy is a tuple $\pi=(\Gamma,m_{0},\delta,\gamma)$ where:
  • $\Gamma$: finite tape alphabet,
  • $m_{0}$: initial tape configuration and head position,
  • $\delta:\Gamma\times Q\times A\rightarrow\Gamma\times\{L,R\}$: tape update and head move function depending only on the scanned symbol,
  • $\gamma:\Gamma\to Dist(C)$: probabilistic control output function depending only on the current scanned tape symbol.
Definition 4.13.
Let $\pi$ be a probabilistic policy and $U$ a controlled probabilistic automaton. The probability of a run $\rho=(q_{0},a_{0})(q_{1}a_{1})\ldots(q_{n-1},a_{n-1})q_{n}$ under $\pi$ is:
$\Pr_{\pi}(\rho)=\prod_{i=0}^{n-1}\left(\sum_{c\in C}\pi(h_{i+1})(c)\cdot P(q_{i},a_{i},c,q_{i+1})\right)$
where $h_{i}=(q_{0},a_{0})\ldots(q_{i-1},a_{i-1})$ is the history up to step $i$.
Probabilistic Languages on Finite Words
In the remainder of this section, to further consider the aspect of formal language recognition by controlled probabilistic automata, let us assume that a controlled automaton has a finite set of states and an additional set of states, F, referred to as the set of accepting states.
Definition 4.14.
Let $U=(Q,A,C,P,Q_{0},F)$ be a controlled probabilistic automaton with a set of accepting states ($F\subseteq Q$). For $w=a_{0}a_{1}\cdots a_{n-1}\in A^{*}$, let $\mathrm{AR}_{\pi}(w)$ denote the set of all accepting runs, if any, generating a given action sequence (word) $w$ under policy $\pi$ in $U$, i.e.,
[…]
Given a threshold $\theta\in[0,1]$, the probabilistic language accepted under a probabilistic policy $\pi$ by $U$ is defined as:
$L^{\theta}_{\pi}(U)=\left\{\,w\in A^{*}\ \middle|\ \sum_{\rho\in\mathrm{AR}_{\pi}(w)}\Pr_{\pi}(\rho)\;\;\geq\theta\right\}.$
where $\Pr_{\pi}(\rho)$ is defined as in Definition 4.13 above.
Proposition 4.15.
Let $U$ be a probabilistic automaton. We define the following classes of languages:
$L^{\theta}_{0}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{ is a memoryless deterministic policy}\},$
$L^{\theta}_{F}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{ is a finite-memory deterministic policy}\},$
$L^{\theta}_{\text{Stack}}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{ is a stack-augmented deterministic policy}\},$
$L^{\theta}_{\text{Tape}}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{ is a tape-augmented deterministic policy}\},$
$L^{\theta}_{H}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{ is a history-dependent deterministic policy}\},$
$L^{\theta}_{0,P}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{is a memoryless probabilistic policy}\},$
$L^{\theta}_{F,P}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{is a finite-memory probabilistic policy}\},$
$L^{\theta}_{\text{Stack},P}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{is a stack-augmented probabilistic policy}\},$
$L^{\theta}_{\text{Tape},P}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{is a tape-augmented probabilistic policy}\},$
$L^{\theta}_{H,P}(U)$ $=\{L^{\theta}_{\pi}(U)\mid\pi\text{is a history-dependent probabilistic policy}\}.$
We have:
$L^{\theta}_{0}(U)\subset L^{\theta}_{F}(U)\subset L^{\theta}_{\text{Stack}}(U)\subset L^{\theta}_{\text{Tape}}(U)\subset L^{\theta}_{H}(U),$
$L^{\theta}_{0,P}(U)\subset L^{\theta}_{F,P}(U)\subset L^{\theta}_{\text{Stack},P}(U)\subset L^{\theta}_{\text{Tape},P}(U)\subset L^{\theta}_{H,P}(U).$
Definition 4.16.
Let $U=(Q,A,C,P,Q_{0},F)$ and $U^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},P^{\prime},Q^{\prime}_{0},F^{\prime})$ be two controlled probabilistic automata with accepting sets $F$ and $F^{\prime}$, respectively, and the same activity alphabet and set of control actions (i.e., $A=A^{\prime}$ and $C=C^{\prime}$). Let $(Q,A,C,P,Q_{0})$ and $(Q^{\prime},A^{\prime},C^{\prime},P^{\prime},Q^{\prime}_{0})$ be two equivalent controlled probabilistic automata under a bisimulation $\gamma$, as in Definition 4.3. Then, $U$ and $U^{\prime}$ are said to be equivalent if:
  • for any $q\in F$, there exists $q^{\prime}\in F^{\prime}$ where $(q,q^{\prime})\in\gamma$; also for any $q^{\prime}\in F^{\prime}$, there exists $q\in Q$ where $(q^{\prime\prime},q)\in\gamma$.
Proposition 4.17.
Let $U$ and $U^{\prime}$ be two equivalent controlled probabilistic automata as in Definition 4.16. Consider the classes of languages as defined in Proposition 4.15. Then, we have:
$L^{\theta}_{0}(U)$ $=L^{\theta}_{0}(U^{\prime}),$
$L^{\theta}_{F}(U)$ $=L^{\theta}_{F}(U^{\prime}),$
$L^{\theta}_{\text{Stack}}(U)$ $=L^{\theta}_{\text{Stack}}(U^{\prime}),$
$L^{\theta}_{\text{Tape}}(U)$ $=L^{\theta}_{\text{Tape}}(U^{\prime}),$
$L^{\theta}_{H}(U)$ $=L^{\theta}_{H}(U^{\prime}),$
$L^{\theta}_{0,P}(U)$ $=L^{\theta}_{0,P}(U^{\prime}),$
$L^{\theta}_{F,P}(U)$ $=L^{\theta}_{F,P}(U^{\prime}),$
$L^{\theta}_{\text{Stack},P}(U)$ $=L^{\theta}_{\text{Stack},P}(U^{\prime}).$
$L^{\theta}_{\text{Tape},P}(U)$ $=L^{\theta}_{\text{Tape},P}(U^{\prime}).$
$L^{\theta}_{H,P}(U)$ $=L^{\theta}_{H,P}(U^{\prime}).$
Also, for equivalent polices $\pi$ and $\pi^{\prime}$ of $U$ and $U^{\prime}$, respectively, as in Definition 4.9, we have:
$L^{\theta}_{\pi,0}(U)$ $=L^{\theta}_{\pi^{\prime},0}(U^{\prime}),$
$L^{\theta}_{\pi,F}(U)$ $=L^{\theta}_{\pi^{\prime},F}(U^{\prime}),$
$L^{\theta}_{\pi,\text{Stack}}(U)$ $=L^{\theta}_{\pi^{\prime},\text{Stack}}(U^{\prime}),$
$L^{\theta}_{\pi,\text{Tape}}(U)$ $=L^{\theta}_{\pi^{\prime},\text{Tape}}(U^{\prime}),$
$L^{\theta}_{\pi,H}(U)$ $=L^{\theta}_{\pi^{\prime},H}(U^{\prime}),$
$L^{\theta}_{\pi,0,P}(U)$ $=L^{\theta}_{\pi^{\prime},0,P}(U^{\prime}),$
$L^{\theta}_{\pi,F,P}(U)$ $=L^{\theta}_{\pi^{\prime},F,P}(U^{\prime}),$
$L^{\theta}_{\pi,\text{Stack},P}(U)$ $=L^{\theta}_{\pi^{\prime},\text{Stack},P}(U^{\prime}).$
$L^{\theta}_{\pi,\text{Tape},P}(U)$ $=L^{\theta}_{\pi^{\prime},\text{Tape},P}(U^{\prime}).$
$L^{\theta}_{\pi,H,P}(U)$ $=L^{\theta}_{\pi^{\prime},H,P}(U^{\prime}).$
Global Language Families for Controlled Probabilistic Automata on Finite Words
We now introduce the global language families associated with controlled probabilistic automata as formalized in Definition 4.14. We define sets of language classes indexed by probabilistic systems and stratified by policy type.
Definition 4.18.
Let $\mathcal{U}$ denote the set of all controlled probabilistic automata (as in Definition 4.2). For each policy class $\Pi\in\{0,F,\text{Stack},\text{Tape},H\}$, define:
$\mathbb{L}^{\theta}_{\Pi,P}:=\{L^{\theta}_{\Pi,P}(U)\mid U\in\mathcal{U}\}\subseteq\mathcal{P}(2^{A^{*}})$
where $L^{\theta}_{\Pi,P}(U):=\{L^{\theta}_{\pi,P}(U)\mid\pi\in\Pi\}$ is the family of languages on finite words of a controlled probabilistic automaton $U$ accepted with probability above a fixed threshold $\theta\in(0,1]$ under a probabilistic policy $\pi$.
Hierarchy
Theorem 4.19.
As in the deterministic setting, we obtain a strict inclusion hierarchy:
$\mathbb{L}^{\theta}_{0,P}\subset\mathbb{L}^{\theta}_{F,P}\subset\mathbb{L}^{\theta}_{\text{Stack},P}\subset\mathbb{L}^{\theta}_{\text{Tape},P}\subset\mathbb{L}^{\theta}_{H,P}.$
Proof Sketch.
Analogous to the proof sketch of Theorem 3.29. ∎
Theorem 4.20.
As in the deterministic setting, we obtain a strict inclusion hierarchy:
$\mathbb{L}^{\theta}_{0,P}\subset\mathbb{L}^{\theta}_{F,P}\subset\mathbb{L}^{\theta}_{\text{Stack},P}\subset\mathbb{L}^{\theta}_{\text{Tape},P}\subset\mathbb{L}^{\theta}_{H,P}.$
Proof Sketch.
We only outline why the inclusions hold and are strict; the argument is a probabilistic lifting of the deterministic hierarchy in Theorem 3.29.
Inclusions. For any controlled probabilistic automaton $U$:
  • A memoryless policy is a special case of a finite-memory policy, so $L^{\theta}_{0,P}(U)\subseteq L^{\theta}_{F,P}(U)$.
  • A finite-memory policy can be encoded by a pushdown controller with bounded stack height, giving $L^{\theta}_{F,P}(U)\subseteq L^{\theta}_{\mathrm{Stack},P}(U)$.
  • A Turing-power controller can simulate a pushdown controller, yielding $L^{\theta}_{\mathrm{Stack},P}(U)\subseteq L^{\theta}_{\mathrm{Tape},P}(U)$.
  • Finally, Turing-power controllers are a special case of general history-dependent policies, so $L^{\theta}_{\mathrm{Tape},P}(U)\subseteq L^{\theta}_{H,P}(U)$.
Taking unions over all $U$ gives the stated chain of inclusions for the global families.
Strictness. For each step, we reuse the deterministic witnesses from Theorem 3.29, viewed as degenerate probabilistic systems whose transition probabilities are 0 or $1$:
  • Lemma 4.21 provides a single probabilistic plant $U$ and a countable family of finite-memory policies $\{\pi_{k}\}$ with distinct regular languages, showing $\mathbb{L}^{\theta}_{0,P}\subset\mathbb{L}^{\theta}_{F,P}$.
  • The pushdown vs. finite-memory and Turing-power vs. pushdown separations are obtained by taking the deterministic plants and controllers from Chapter 3 and interpreting their transitions as probabilistic with probability $1$; the acceptance probabilities are still 0 or $1$, so the same separating languages appear in $\mathbb{L}^{\theta}_{\text{Stack},P}$ and $\mathbb{L}^{\theta}_{\text{Tape},P}$ but not in the smaller classes.
  • Finally, the history-dependent vs. Turing-power separation is established by the same kind of non-realizable language family as in Theorem 3.29, again interpreted in the probabilistic setting via transitions of probability $1$.
Thus, each inclusion in the chain is proper, yielding the hierarchy. ∎
The following lemma provides a constructive witness for the strictness claimed in Theorem 4.20.
Lemma 4.21.
There exists a controlled probabilistic automaton $U$ such that $L_{F}^{\theta}(U)$ contains a countably infinite family of distinct regular languages.
Proof Sketch.
Consider a plant $U$ with a single probabilistic state $q$ and self-loop on $a$ for every control. For each $k\geq 1$, define a finite-memory policy $\pi_{k}$ with a $k$-state deterministic counter controlling acceptance: the plant is absorbing, and acceptance is determined entirely by the controller.
The acceptance probability under $\pi_{k}$ is $1$ iff the input belongs to $(a^{k})^{*}$, and 0 otherwise. Since probabilistic policies subsume deterministic ones, $(a^{k})^{*}\in L_{F}^{\theta}(U)$ for all $k$, and the resulting family is countably infinite. ∎
Remark 4.22.
Lemma 4.21 is a probabilistic analog of Lemma 3.30 for deterministic control.
Emptiness Problem for Controlled Probabilistic Automata
Definition 4.23.
Let $U=(Q,A,C,P,Q_{0},F)$ be a controlled probabilistic automaton (CPA), let $\Pi$ be a class of admissible policies, and let $\theta\in[0,1]$ be a probability threshold. The emptiness problem for CPAs is to decide whether
$\exists\,\pi\in\Pi\ \text{such that}\ L_{\theta}^{\pi}(U)\neq\emptyset,$
equivalently,
$\exists\,\pi\in\Pi,\ \exists\,w\in A^{*}\ \text{such that}\ \Pr_{\pi}(\mathrm{AR}(w))\geq\theta,$
where $\mathrm{AR}(w)$ denotes the set of accepting runs, if any, that generate a given finite action sequence (word) w under policy $\pi$ in $U$ (see definition 4.14).
Proposition 4.24.
The emptiness problem for controlled probabilistic automata (CPAs) is undecidable for $\theta>0$ under general (e.g., history-dependent) policies.
Proof Sketch.
The undecidability of the emptiness problem for probabilistic automata with a non-strict cutpoint was first established by Rabin [Rabin1963] through a reduction from the Post correspondence problem. Paz [Paz1971] provides a comprehensive exposition of this result and related decision problems, including the structure of probabilistic languages and the behavior of cutpoints. Madani, Hanks, and Condon [madani2003] later extended these ideas to probabilistic planning, showing that undecidability persists in control-oriented settings. Together, these results establish that no algorithm can decide emptiness for probabilistic automata under a non-strict cutpoint semantics.
Since a controlled probabilistic automaton with a singleton control set and a trivial policy behaves identically to a probabilistic finite automaton, the emptiness problem for CPAs reduces directly to Rabin’s classical result, proving undecidability. ∎
Remark 4.25.
For strict or isolated cutpoints, certain decision problems for probabilistic automata become decidable; however, these restrictions are not considered in this work.
Probabilistic Languages on Infinite Words
Finally, we consider the reactive aspect of a controlled probabilistic automaton by considering its behavior under an infinite sequence of activity completions as follows.
Definition 4.26.
A controlled probabilistic Büchi automaton is a 6-tuple $U_{\omega}=(Q,A,C,P,Q_{0},F)$ where:
  • $(Q,A,C,P,Q_{0})$ is a controlled probabilstic automaton,
  • $F\subseteq Q$ is the set of accepting states (visited infinitely often).
Definition 4.27.
Let $U_{\omega}=(Q,A,C,P,Q_{0},F)$ be a controlled probabilistic Büchi automaton, and let $\pi:(Q\times A)^{*}\to\text{Dist}(C)$ be a probabilistic policy. We can construct a probability space $(\Omega,\mathcal{F},\mathbb{P}_{\pi})$ under policiy $\pi$ as follows. Define the sample space of infinite runs as $\Omega=(Q\times A)^{\omega}$, and let $\mathcal{F}$ be the $\sigma$-algebra generated by cylinder sets [billingsley1995probability] of the form:
$\left\{(Q\times A)^{n}\times Q\;\middle|\;n>0\right\}.$
For any finite run
$\rho[0..n]=(q_{0},a_{0})(q_{1},a_{1})\cdots(q_{n-1},a_{n-1})q_{n},$
define the probability of this finite run under policy $\pi$ as:
$\mathbb{P}_{\pi}(\rho[0..n])=\prod_{i=0}^{n-1}\left(\sum_{c\in C}\pi(h_{i+1})(c)\cdot P(q_{i},a_{i},c,q_{i+1})\right),$
where $h_{i}=(q_{0},a_{0})(q_{1},a_{1})\cdots(q_{i-1},a_{i-1})$ is the prefix history leading up to time step $i$. By Kolmogorov’s extension theorem [Kolmogorov1933, billingsley1995probability], this defines a unique probability measure $\mathbb{P}_{\pi}$ on $(\Omega,\mathcal{F})$.
Definition 4.28.
Let $U_{\omega}=(Q,A,C,P,Q_{0},F)$ be a controlled probabilistic Büchi automaton with a set of accepting states ($F\subseteq Q$). For $w=a_{0}a_{1}a_{2}\cdots\in A^{\omega}$, let $\mathrm{AIR}_{\pi}(w)$ denote the set of all accepting infinite runs, if any, generating a given action sequence $w$ under policy $\pi$ in $U_{\omega}$, i.e.,
[…]
where $\text{Inf}(\rho)$ is the set of states that appear infinitely often in the projection of $\rho$ onto $Q$. Given a threshold $\theta\in[0,1]$, the probabilistic $\omega$-language accepted under a probabilistic policy $\pi$ by $U_{\omega}$ is defined as:
$L^{\omega,\theta}_{\pi}(U_{\omega})=\left\{\,w\in A^{\omega}\ \middle|\ \sum_{\rho\in\mathrm{AIR}_{\pi}(w)}\mathbb{P}_{\pi}(\rho)\;\;\geq\theta\right\}.$
Proposition 4.29.
Let $U_{\omega}$ be a controlled probabilistic Buchi automaton. We define the following classes of probabilistic $\omega$-languages:
$L^{\omega,\theta}_{0}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a memoryless deterministic policy}\},$
$L^{\omega,\theta}_{F}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a finite-memory deterministic policy}\},$
$L^{\omega,\theta}_{\text{Stack}}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a stack-augmanted deterministic policy}\},$
$L^{\omega,\theta}_{\text{Tape}}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a tape-augemted deterministic policy}\},$
$L^{\omega,\theta}_{H}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a history-dependent deterministic policy}\},$
$L^{\omega,\theta}_{0,P}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a memoryless probabilistic policy}\},$
$L^{\omega,\theta}_{F,P}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a finite-memory probabilistic policy}\},$
$L^{\omega,\theta}_{\text{Stack,P}}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a stack-augmanted probabilistic policy}\},$
$L^{\omega,\theta}_{\text{Tape,P}}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{ is a tape-augemted probabilistic policy}\},$
$L^{\omega,\theta}_{H,P}(U_{\omega})$ $=\{L^{\omega,\theta}_{\pi}(U_{\omega})\mid\pi\text{is a history-dependent probabilistic policy}\}.$
We have:
$L^{\omega,\theta}_{0}(U_{\omega})\subset L^{\omega,\theta}_{F}(U_{\omega})\subset L^{\omega,\theta}_{\text{Stack}}(U_{\omega}\subset L^{\omega,\theta}_{\text{Tape}}(U_{\omega})\subset L^{\omega,\theta}_{H}(U_{\omega}),$
$L^{\omega,\theta}_{0,P}(U_{\omega})\subset L^{\omega,\theta}_{F,P}(U_{\omega})\subset L^{\omega,\theta}_{\text{Stack,P}}(U_{\omega})\subset L^{\omega,\theta}_{\text{Tape,P}}(U_{\omega})\subset L^{\omega,\theta}_{H,P}(U_{\omega}).$
Definition 4.30.
Let $U_{\omega}=(Q,A,C,P,Q_{0},F)$ and $U_{\omega}^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},P^{\prime},Q^{\prime}_{0},F^{\prime})$ be two controlled probabilistic Buchi automata with the same activity alphabet and set of control actions (i.e., $A=A^{\prime}$ and $C=C^{\prime}$). Let $(Q,A,C,P,Q_{0})$ and $(Q^{\prime},A^{\prime},C^{\prime},P^{\prime},Q^{\prime}_{0})$ be two equivalent controlled probabilistic automata under a bisimulation $\gamma$, as in Definition 4.3. Then, $U_{\omega}$ and $U^{\prime}_{\omega}$ are said to be equivalent if: be the sets of accepting states for $\mathcal{U}_{\omega}$ and $\mathcal{U}_{\omega}^{\prime}$, respectively, such that
  • for any $q\in F$, there exists $q^{\prime}\in F^{\prime}$ where $(q,q^{\prime})\in\gamma$; also for any $q^{\prime}\in F^{\prime}$, there exists $q\in Q$ where $(q^{\prime\prime},q)\in\gamma$.
Proposition 4.31.
Let $U_{\omega}$ and $U^{\prime}_{\omega}$ be two equivalent controlled probabilistic Buchi automata as in Definition 4.30. Consider the classes of probabilistic $\omega$-languages as defined in Proposition 4.29. Then, we have:
$L^{\omega,\theta}_{0}(U_{\omega})$ $=L^{\omega,\theta}_{0}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{F}(U_{\omega})$ $=L^{\omega,\theta}_{F}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\text{Stack}}(U_{\omega})$ $=L^{\omega,\theta}_{H}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\text{Tape}}(U_{\omega})$ $=L^{\omega,\theta}_{H}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{H}(U_{\omega})$ $=L^{\omega,\theta}_{H}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{0,P}(U_{\omega})$ $=L^{\omega,\theta}_{0,P}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{F,P}(U_{\omega})$ $=L^{\omega,\theta}_{F,P}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\text{Stack},P}(U_{\omega})$ $=L^{\omega,\theta}_{H,P}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\text{Tape},P}(U_{\omega})$ $=L^{\omega,\theta}_{H,P}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{H,P}(U_{\omega})$ $=L^{\omega,\theta}_{H,P}(U_{\omega}^{\prime}).$
Also, for equivalent polices $\pi$ and $\pi^{\prime}$ of $U_{\omega}$ and $U_{\omega}^{\prime}$, respectively, as in Definition 4.9, we have:
$L^{\omega,\theta}_{\pi,0}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},0}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,F}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},F}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,\text{Stack}}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},H}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,\text{Tape}}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},H}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,H}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},H}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,0,P}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},0,P}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,F,P}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},F,P}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,\text{Stack},P}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},H,P}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,\text{Tape},P}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},H,P}(U_{\omega}^{\prime}),$
$L^{\omega,\theta}_{\pi,H,P}(U_{\omega})$ $=L^{\omega,\theta}_{\pi^{\prime},H,P}(U_{\omega}^{\prime}).$
When $\theta=0^{+}$ or $\theta=1$, where $0^{+}$ denotes a sufficiently small positive real number, the model is said to have a positive acceptance or almost-sure acceptance semantics, respectively [baier2012probabilistic].
We now consider the global language families associated with controlled probabilistic Büchi automata as formalized in Definition  4.28. We define sets of language classes indexed by controlled probabilistic systems and stratified by policy type.
Global Language Families for Controlled Probabilistic Büchi Automata
Definition 4.32.
Let $\mathcal{U}_{\omega}$ be the set of all probabilistic controlled Büchi automata (Definition  4.26). For each policy class $\Pi\in\{0,F,\text{Stack},\text{Tape},H\}$, define:
$\mathbb{L}^{\omega,\theta}_{\Pi,P}:=\left\{L^{\omega,\theta}_{\Pi}(U_{\omega})\mid U_{\omega}\in\mathcal{U}_{\omega}\right\}\subseteq\mathcal{P}(2^{A^{\omega}})$
where $L^{\omega,\theta}_{\Pi,P}(U_{\omega}):=\{L^{\omega}_{\pi}(U_{\omega})\mid\pi\in\Pi\}$ is the class of languages of infinite words accepted with probability at least $\theta>0$ under policy $\pi$.
Semantics
Each $L^{\omega,\theta}_{\pi}(U_{\omega})$ contains infinite words that lead to accepting runs with probability exceeding the threshold, satisfying the Büchi condition under probabilistic execution.
Hierarchy
Theorem 4.33.
There is again a strict containment structure:
$\mathbb{L}^{\omega,\theta}_{0,P}\subset\mathbb{L}^{\omega,\theta}_{F,P}\subset\mathbb{L}^{\omega,\theta}_{\text{Stack},P}\subset\mathbb{L}^{\omega,\theta}_{\text{Tape},P}\subset\mathbb{L}^{\omega,\theta}_{H,P}.$
Proof Sketch.
Analogous to the proof sketch of Theorem 3.29. ∎
Proof Sketch.
The argument parallels Theorem 3.29, now lifted to the probabilistic Büchi setting. Each inclusion follows because stronger controller classes strictly extend weaker ones: memoryless $\subset$ finite-memory $\subset$ stack-augmented $\subset$ tape-based $\subset$ history-dependent. Strictness is witnessed by adapting the corresponding deterministic constructions of Theorem 3.29, which remains valid under degenerate probabilistic transitions (probability 0 or $1$). Thus, each containment in the stated chain is proper. ∎
Lemma 4.34.
There exists a controlled probabilistic Büchi automaton $U_{\omega}$ such that $L_{F,P}^{\omega,\theta}(U_{\omega})$ contains a countably infinite family of distinct $\omega$-regular languages.
Proof Sketch.
This is the probabilistic Büchi analogue of Lemma 3.52. We use the same two-state plant with accepting state $q_{\mathrm{ok}}$ and rejecting sink $q_{\mathrm{sink}}$, and the family of finite-memory controllers $\{\pi_{k}\}_{k\geq 1}$ that count the number of consecutive $a$’s before each $b$ modulo $k$. Each controller $\pi_{k}$ enforces the $\omega$-regular language $(a^{k}b)^{\omega}$ under the almost-sure semantics ($\theta=1$), since runs remain in $q_{\mathrm{ok}}$ precisely when every block of $a$’s has length $k$. Because transitions may be taken as degenerate (probability 1), the probabilistic interpretation does not change the behavior. Hence $L_{F,P}^{\omega,\theta}(U_{\omega})$ contains the countably infinite family $\{(a^{k}b)^{\omega}\mid k\geq 1\}$. ∎
Emptiness Problem for Controlled Probabilistic Büchi Automata
Definition 4.35.
Let $U_{\omega}=(Q,A,C,P,Q_{0},F)$ be a controlled probabilistic Büchi automaton, let $\theta\in(0,1]$ be a fixed threshold, and let $\Pi$ denote a class of admissible policies. The emptiness problem asks whether there exists a policy $\pi\in\Pi$ and an infinite word $w=a_{0}a_{1}a_{2}\cdots\in A^{\omega}$ such that the set of accepting runs over $w$ under $\pi$ has probability at least $\theta$; formally,
$\exists\,\pi\in\Pi,\ \exists\,w\in A^{\omega}\ \text{such that}\ \mathbb{P}_{\pi}(\mathrm{AR}(w))\geq\theta,$
where $\mathrm{AR}(w)$ denotes the set of accepting runs labeled by $w$ as in Definition 4.27. Equivalently, the problem asks whether the threshold language $L_{\omega,\theta}^{\pi}(U_{\omega})$ is nonempty for some $\pi\in\Pi$, i.e.,
$\exists\,\pi\in\Pi\ \text{such that}\ L_{\omega,\theta}^{\pi}(U_{\omega})\neq\emptyset.$
Proposition 4.36.
The emptiness problem for controlled probabilistic Büchi automata (Definition 4.35) is undecidable in general. That is, given a CPBA $U_{\omega}=(Q,A,C,P,Q_{0},F)$, a threshold $\theta\in(0,1]$, and a class of admissible policies $\Pi$, no algorithm can decide whether
$\exists\,\pi\in\Pi\ \text{such that}\ L_{\omega,\theta}^{\pi}(U_{\omega})\neq\emptyset,$
equivalently, $\exists\,\pi\in\Pi,\ \exists\,w\in A^{\omega}$ with $\Pr_{\pi}(\mathrm{AR}(w))\geq\theta.$ The result holds even for deterministic history-dependent policies.
Proof Sketch.
The undecidability of the emptiness problem for probabilistic automata with a non-strict cutpoint was established by Rabin [Rabin1963] for probabilistic finite automata. This result extends to probabilistic Büchi automata (PBAs) under positive semantics, as shown in [baier2012probabilistic, courcoubetis1995complexity]. Specifically, given a PBA $A=(Q,A,\delta,q_{0},F)$ and a threshold $\theta\in(0,1]$, deciding whether there exists $w\in A^{\omega}$ with $\Pr(A\text{ accepts }w)\geq\theta$ is undecidable, even for $\theta=0$.
To lift this to controlled probabilistic Büchi automata (CPBAs), we construct $U_{\omega}=(Q,A,C,P,Q_{0},F)$ with $C=\{c_{0}\}$ and $P(q,a,c_{0},q^{\prime})=\delta(q,a,q^{\prime})$. A trivial policy $\pi$ that always selects $c_{0}$ ensures that $U_{\omega}$ under $\pi$ has the same accepting behavior as $A$.
Hence,
$L^{\pi}_{\omega,\theta}(U_{\omega})\neq\emptyset\quad\Leftrightarrow\quad L_{\omega,\theta}(A)\neq\emptyset.$
Since the latter problem is undecidable for PBAs under positive semantics, the emptiness problem for CPBAs is also undecidable. The result continues to hold for deterministic history-dependent policies and fixed thresholds. ∎
Remark 4.37 (Positive vs. Almost-Sure Semantics).
The undecidability of the emptiness problem stated above holds for positive semantics, where a word is accepted if its acceptance probability is strictly greater than a threshold $\theta$. In contrast, under almost-sure semantics, where acceptance requires probability $1$, several decision problems become decidable for important subclasses of probabilistic automata [baier2012probabilistic, courcoubetis1995complexity]. This sharp contrast highlights the subtle role of cutpoints in probabilistic verification.
These results highlight the sharp contrast between deterministic and probabilistic settings: while emptiness is tractable for regular and pushdown systems, the introduction of probabilistic thresholds leads quickly to undecidability.
To analyze decision-making in controlled probabilistic systems, it is essential to extend the structural semantics of controlled probabilistic automata with quantitative measures. The reward model introduced below provides such a foundation by associating costs or benefits with both control decisions and activity completions. This framework allows us to formalize performance and optimality criteria, preparing the ground for policy synthesis and complexity analyses in the discrete-time and continuous-time settings explored later in this section.
Reward Model
Definition 4.38.
Let $U=(Q,A,C,P,Q_{0})$ be a controlled probabilistic automaton. We can define a reward structure $r=(r^{\prime},r^{\prime\prime})$, where $r^{\prime}$ is a bounded Borel measurable function
$r^{\prime}:Q\times C\to\mathbb{R}_{+},$
referred to as the reward rate function, and $r^{\prime\prime}$ is a bounded Borel measurable function
$r^{\prime\prime}:Q\times A\times C\times Q\to\mathbb{R}_{+},$
called the activity reward function.
The model can be viewed to evolve over discrete time, where the completion of an activity in a state triggers a control decision that determines the next state. Let $X(i)$, $Y(i)$, and $Z(i)$ represent the state, control action, and activity completed at time $i$, respectively, for $i\in\mathbb{N}_{+}$. The reward incurred during a period $k^{\prime},k]$, for $k^{\prime},k\in\mathbb{N}_{+}$ with $k^{\prime}<k$, is defined as the random variable
$R=\sum_{i=k^{\prime}}^{k-1}r^{\prime}(X(i),Y(i))+\sum_{i=k^{\prime}}^{k-1}r^{\prime\prime}(X(i),Z(i),Y(i),X(i+1)).$ (4.2)
The random variable $R$ is well-defined under the measurability and boundedness conditions above. To compare rewards, we use the stochastic ordering $\leq_{st}$ on random variables: for $X,Y$, we write $X\leq_{st}Y$ if and only if $\Pr[X>x]\leq\Pr[Y>x]$ for all $x\in\mathbb{R}_{+}$.
Let $U=(Q,A,C,P,Q_{0})$ be a controlled probabilistic automaton (CPA) with reward structure $r$, and let $\mathcal{P}$ denote a class of admissible policies. For any $k^{\prime},k\in\mathbb{N}_{+}$ with $k^{\prime}\leq k$, $\pi\in\mathcal{P}$, and $q\in Q$, define
$R_{k^{\prime}}^{k}(\pi|q)\equiv\text{the reward accrued during }[k^{\prime},k]\text{ given }X(k^{\prime})=q.$
A policy $\pi^{*}\in\mathcal{P}$ is called an optimal policy for $(V,r)$ over $k^{\prime},k]$ if for all $\pi\in\mathcal{P}$ and $q\in Q$,
$R_{k^{\prime}}^{k}(\pi|q)\leq_{st}R_{k^{\prime}}^{k}(\pi^{*}|q).$
However, such an optimal policy is quite strong and may not always exist. Accordingly, we may consider other relaxed versions of the notion of optimality as follows.
Let $R_{k^{\prime}}^{k}(\pi\mid q_{0})$ denote the random variable representing the total accumulated reward over the time interval $k^{\prime},k]$, when starting from state $q_{0}$ and following policy $\pi$. We assume that rewards are bounded and measurable, and that the underlying stochastic process is non-explosive, so that the expected total reward is well defined for all admissible policies.
Formally, for a run $\rho=(q_{0},a_{0})(q_{1},a_{1})\cdots(q_{n-1},a_{n-1})q_{n}$ generated under $\pi$, the total reward is computed as the integral of the reward rate associated with the visited states and control actions over time, as in Equation 4.2. For expected-total and discounted-reward objectives, the controller’s goal is to select a policy that maximizes $\mathbb{E}[R_{k^{\prime}}^{k}(\pi\mid q_{0})]$ or its discounted analogue.
For models with finite state and control action spaces, the existence of stationary deterministic optimal policies is well established for expected-total and discounted-reward objectives. For time-bounded objectives, optimal policies may require history dependence. In more general settings (e.g., infinite state spaces or general timing distributions), the existence of exact optimal policies depends on regularity assumptions such as boundedness, continuity, and compactness of action spaces. Even when these conditions do not hold, $\varepsilon$-optimal policies can be obtained under broad conditions. These results will be further elaborated in more detail in this section, as follows, while providing the mathematical basis for the reward-based formulation of control problems in the following section.
While controlled probabilistic automata offer a general representation of decision-driven dynamics, many theoretical and algorithmic results are most naturally expressed in the language of discrete-time Markov decision processes (DTMDPs). The following definition specializes this classical framework to our setting by incorporating the activity set and explicit control structure, ensuring consistency with the preceding reward semantics and policy definitions.
Discrete-Time Markov Decision Process (DTMDP)
We now introduce a popular and general framework commonly used in most theoretical decision-making problems that evolve over discrete time, and its relation to our proposed models is as follows.
Definition 4.39.
A discrete-time Markov decision process (DTMDP) is defined as a 5-tuple $D=(Q,C,Q_{0},P)$, where:
  • Q is a countable set of states,
  • C is a finite set of control actions,
  • $Q_{0}$ is the initial state distribution where $\sum_{q\in Q}Q_{0}(q)=1$,
  • $P:Q\times C\times Q\to[0,1]$ is the transition probability function, where for any $q\in Q$ and $c\in C$,
    $\sum_{q^{\prime}\in Q}P(q,c,q^{\prime})=1.$
Definition 4.40.
Let $U=(Q,A,C,P,Q_{0})$ be a controlled probabilistic automaton (CPA). Then, $U$ is said to be represented by a discrete-time Markov decision process (DTMDP) $D=(Q,C,Q_{0},P^{\prime})$ defined as:
$P^{\prime}(q,c,q^{\prime})=\sum_{a\in A}P(q,a,c,q^{\prime}).$
In the above model representation, the activity component becomes implicit, and the DTMDP models the same transition dynamics by summarizing the rate and probability information from the CPA. Thus, a DTMDP model can be viewed as an abstract representation of a CPA model. More formally, we have:
Proposition 4.41.
Let $U=(Q,A,C,P,Q_{0})$ and $D=(Q,C,Q_{0},\lambda,P^{\prime})$ be a controlled probabilistic automaton (CPA) and a discrete-time Markov decision process (DTMDP), respectively, as defined in Definition  4.40 above. Then, there exist policies $\pi_{U}$ and $\pi_{D}$ in $U$ and $D$, respectively, such that the state processes of $U$ and $D$ under policies $\pi_{U}$ and $\pi_{D}$ will be isomorphic respectively.
With the reward and policy structures in place, we now turn to the computational aspects of policy synthesis in finite-state DTMDPs. Understanding the complexity of finding optimal strategies is crucial for evaluating the tractability of controller synthesis in our framework, and it directly parallels the results known for continuous-time models derived from Controlled SANs as discussed in the following section.
Optimal Policies for Finite-State DTMDPs
For a controlled probabilistic automaton $U=(Q,A,C,P,Q_{0})$ with finite state space $Q$ and control action set $C$, the induced discrete-time Markov decision process (DTMDP) admits standard algorithmic solutions for synthesizing optimal policies under bounded reward structures. The computational complexity depends on the quantitative objective, as illustrated by the following property:
Proposition 4.42.
Let $D$ be a finite-state discrete-time Markov decision process with bounded rewards and a finite control action set. Then:
  1. 1.
    An optimal stationary deterministic policy exists for both expected total reward and discounted reward objectives.
  2. 2.
    Such a policy can be computed in polynomial time.
  3. 3.
    For time-bounded reachability or reward objectives, the optimal value can be computed in PSPACE, and the problem is PSPACE-complete in general.
  4. 4.
    In the time-bounded case, optimal policies may require history dependence.
Proof Sketch.
The existence of optimal stationary deterministic policies for expected total and discounted reward objectives follows from classical results on finite-state MDPs [Howard71, puterman1994markov]. Intuitively, dynamic programming or linear programming methods can compute an optimal policy in polynomial time since the state and action spaces are finite. For time-bounded reachability or reward objectives, PSPACE-completeness follows from standard reductions used in model checking of MDPs [courcoubetis1995complexity], and history-dependent policies may be required to achieve optimality. ∎
Remark 4.43.
For time-bounded objectives, optimality may require history-dependent policies because the choice of action can depend on the remaining time horizon, which cannot be encoded by a stationary deterministic policy alone [courcoubetis1995complexity, baier2008principles].
Table 1 summarizes classical results on the existence of optimal policies and their computational characteristics for finite-state DTMDPs. It outlines typical quantitative objectives, the corresponding structure of optimal policies, and the algorithmic methods commonly used to compute them efficiently.
Table 1: Complexity and policy characteristics for optimal control in finite-state DTMDPs, categorized by quantitative objective.
Objective Optimal Policy Type Complexity Class Solution Method
Expected Total Reward Memoryless deterministic Polynomial time Linear programming [Howard71, puterman1994markov]
Discounted Reward Memoryless deterministic Polynomial time Value iteration / policy iteration [Howard71, puterman1994markov]
Finite-Horizon Reachability / Reward May require history dependence PSPACE-complete Dynamic programming over horizon [courcoubetis1995complexity]
These complexity results show that expected total and discounted reward problems are efficiently solvable with simple policy structures, while time-bounded objectives are computationally harder and may require policies with memory. These findings provide the foundation for subsequent analysis of controlled probabilistic automata and their reductions to DTMDPs for policy synthesis.
Although many decision-making problems involve finite-state abstractions, real-world systems often yield DTMDPs with countably infinite state spaces—for example, when queue lengths or resource counters are unbounded. The following discussion outlines the conditions under which stationary or history-dependent optimal policies can be guaranteed, and highlights approximation techniques used when exact solutions are infeasible.
Optimal Policies for Infinite-State DTMDPs
When a controlled probabilistic automaton $U=(Q,A,C,P,Q_{0})$ induces a discrete-time Markov decision process (DTMDP) $D=(Q,C,Q_{0},P^{\prime})$ with a countably infinite state space $Q$, the existence and synthesis of optimal policies present significant challenges beyond those seen in the finite-state case. While finite-state DTMDPs always admit memoryless deterministic optimal policies for expected total and discounted rewards, additional regularity conditions are required to guarantee well-defined solutions in the infinite-state setting. We have the following property:
Proposition 4.44.
Consider a discrete-time Markov decision process $D=(Q,C,Q_{0},P^{\prime})$ induced by a controlled probabilistic automaton with a countably infinite state space $Q$ and bounded, measurable reward function $r:Q\times C\to\mathbb{R}$. Assuming non-explosive dynamics, the following hold:
  1. 1.
    The optimal value function $V^{*}$ is finite and well-defined.
  2. 2.
    There exists an optimal policy $\pi^{*}$ attaining the supremum expected reward.
  3. 3.
    For discounted rewards ($0<\gamma<1$), $V^{*}$ can be obtained via value iteration or policy iteration.
  4. 4.
    Exact computation of $\pi^{*}$ is generally infeasible, but $\varepsilon$-optimal policies can be approximated through methods such as state-space truncation, uniformization, or grid-based abstraction.
  5. 5.
    Under additional continuity and compactness assumptions on the state and action spaces, measurable selectors and stationary $\varepsilon$-optimal policies exist.
Proof Sketch.
These results follow from the classical theory of Markov decision processes over countable and Borel state spaces. Bounded rewards and non-explosiveness guarantee that the Bellman optimality operator is well-defined and that the value function is finite. For discounted reward criteria, the Bellman operator is a contraction mapping, so convergence of value iteration follows from the Banach fixed-point theorem. Existence of optimal stationary deterministic policies under continuity and compactness assumptions is established using measurable selection theorems. Approximation schemes for infinite-state MDPs are standard techniques for constructing $\varepsilon$-optimal policies. For details, see [puterman1994markov, feinberg2002handbook, HernandezLerma1996]. ∎
Remark 4.45.
Unlike the finite-state case, optimal policies in infinite-state DTMDPs may require additional regularity assumptions, and exact computation is often infeasible. Approximation methods are therefore essential in practical settings, particularly for safety-critical or AI-driven control applications.
While exact computation of optimal policies for infinite-state DTMDPs is often intractable, a variety of approximation methods can yield $\varepsilon$-optimal solutions in practice. Table 2 summarizes standard methods for synthesizing optimal or $\varepsilon$-optimal policies for both finite-state and infinite-state DTMDPs. It highlights key objectives, policy existence guarantees, and common approximation techniques used when exact computation is infeasible in infinite-state settings.
Table 2: Comparison of optimal policy properties and solvability for finite- and infinite-state DTMDPs.
Setting / Objective Optimal Policy Type Existence / Complexity Solution Method
Finite-State, Expected Total Reward Memoryless deterministic Polynomial time Linear programming [puterman1994markov]
Finite-State, Discounted Reward Memoryless deterministic Polynomial time Value or policy iteration [puterman1994markov]
Finite-State, Time-Bounded Objectives May require history dependence PSPACE-complete Dynamic programming over horizon [courcoubetis1995complexity]
Infinite-State, Expected or Discounted Reward Stationary deterministic (under continuity) Existence guaranteed; exact computation typically infeasible Value iteration with truncation, uniformization, or grid-based approximation [puterman1994markov]
Infinite-State, Time-Bounded Objectives Often history-dependent or approximate only No general complexity bounds; $\varepsilon$-optimal policies achievable State abstraction and approximation [courcoubetis1995complexity, feinberg2002handbook, HernandezLerma1996]
5 Stochastic Models
In the previous section, nondeterminacy has been treated in a probabilistic manner. We now present models that represent both nondeterminacy and parallelism probabilistically. This is accomplished by assigning certain parameters to timed activities and viewing the model in a stochastic setting.
Model Structure
Definition 5.1.
A controlled stochastic activity network is an 13-tuple $(P,IA,TA,CA,IG,OG,IR,IOR,TOR,IP,\overline{F},\overline{\rho},\overline{\Pi})$ where:
  • ( P, IA, TA, CA, IG, OG, IR, IOR, TOR, IP) is a probabilistic activity network,
  • $\overline{F}=\{\overline{F}(.|\mu,a);\mu\in{\mathcal{N}}^{n},a\in TA\}$ is the set of activity time distribution functions, where $n=|P|$ and, for any $\mu\in{\mathcal{N}}^{n}$ and $a\in TA$, $\overline{F}(.|\mu,a)$ is a probability distribution function,
  • $\overline{\rho}:\;{\mathcal{N}}^{n}\times TA\longrightarrow{\mathcal{R}_{+}}$ is the enabling rate function, where $n$ is defined as before.
  • $\overline{\Pi}:\;{\mathcal{N}}^{n}\times TA\longrightarrow\{true,false\}$ is the reactivation predicate,
Model Behavior
The behavior of the above model is similar to that of a controlled probabilistic activity network, except that here the notion of timing is explicitly considered. When instantaneous activities are enabled, they complete instantaneously. Enabled timed activities, on the other hand, require some time to complete. A timed activity becomes active as soon as it is enabled and remains so until it completes; otherwise, it is inactive. Consider a controlled stochastic activity network $M$ as in Definition  5.1. Suppose, at time $t$, a timed activity completes, and $\mu$ is the stable marking of $M$ immediately after $t$. A timed activity $a$ is activated at $t$, if $a$ is enabled in $\mu$ and one of the following occurs:
  • $a$ is inactive immediately before $t$,
  • $a$ completes at $t$,
  • $\overline{\Pi}(\mu,a)=true$.
Whenever the above happens $a$ is assigned an activity time $\tau$, where $\tau$ is a random variable with a probability distribution function $\overline{F}(.|\mu,a)$. When a timed activity $a$ is enabled in a stable marking $\mu$, it is processed with a rate $\overline{\rho}(\mu,a)$. A timed activity completes whenever it is processed for its activity time. Upon completion of an activity, the next marking occurs immediately.
Semantic Model
The above summarizes the behavior of a controlled stochastic activity network. This behavior may be studied more formally using the following concepts.
Definition 5.2.
A controlled stochastic automaton is a 7–tuple
$(Q,\,A,\,C,\,P,\,Q_{0},\,F,\,\rho,\,\Pi),$
where:
  • $(Q,A,C,P,Q_{0}),$ is a controlled probabilistic automaton,
  • $F=\{F(.|q,a):q\in Q,a\in A\}$ is the set of activity time distribution functions where for any $q\in Q$ and $a\in A$, $F(.|q,a)$ is a probability distribution function,
  • $\rho:Q\times A\longrightarrow{\mathcal{R}_{+}}$ is the enabling rate function,
  • $\Pi:Q\times A\longrightarrow\{false,true\}$ is the reactivation predicate.
Controlled stochastic automata are employed to abstractly model both the nondeterminism and parallelism inherent in distributed real-time systems. In the proposed definition, the first case captures nondeterminism in a probabilistic manner, within the framework of controlled probabilistic automata. Notably, this formulation is closely tied to the control actions that govern system behavior. In contrast, the remaining three cases address the modeling of parallelism by incorporating the system’s real-time characteristics. These timing aspects are explicitly defined to be orthogonal to the control actions, ensuring a modular and compositional representation of timing and control. These abstract models are used as semantic models for controlled stochastic activity networks. They can be viewed as dynamic systems similar to these latter models, with a similar behavior as follows.
Consider a controlled stochastic automaton $V=(Q,A,C,P,Q_{0},F,\rho,\Pi)$ as in the above definition. An activity becomes active as soon as it is enabled and remains so until it completes; otherwise, it is inactive.Suppose, at time $t$, an activity completes, and $q$ is the state of $V$ Immediately after $t$. $a$ is activated at $t$ if one of the following events occurs:
  • $a$ is inactive immediately before $t$,
  • $a$ completes at $t$,
  • $\Pi(q,a)=true$.
Whenever the above happens $a$ is assigned an activity time $\tau$, where $\tau$ is a random variable with a probability distribution function $F(.|q,a)$. When an activity $a$ is enabled in a s state $q$, it is processed with a rate $\rho(q,a)$. An activity completes whenever it is processed for its activity time. Upon completion of an activity, the next marking occurs immediately.
Bisimulation
Next, we define a notion of equivalence for Controlled stochastic automata, which is similar to the one proposed for the uncontrolled model [movaghar2001stochastic].
Definition 5.3.
Two controlled stochastic automata $V=(Q,A,C,P,Q_{0}.F,\rho,\Pi)$ and $V^{\prime}=(Q^{\prime},A^{\prime},C^{\prime},$ $P^{\prime},Q^{\prime}_{0}.F^{\prime},\rho^{\prime},\Pi^{\prime})$ with the same activity alphabet and set of control actions ($A=A^{\prime}$ and $C=C^{\prime}$) are said to be equivalent if:
  • there is a bisimulation $\gamma$ between two controlled probabilistic automata $(Q,A,C,P,Q_{0})$ and $(Q^{\prime},A^{\prime},C^{\prime},P,Q^{\prime}_{0})$,
  • for any $q\in Q$, $q^{\prime}\in Q^{\prime}$ such that $(q,q^{\prime})\in\gamma$ and $a\in A$, we have $F(.|q,a)=F^{\prime}(.|a,q^{\prime})$, $\rho(q,a)=\rho^{\prime}(q^{\prime})$, and $\Pi(q,a)=\Pi^{\prime}(q^{\prime},a)$.
We say that $\gamma$ is also a bisimulation between $V$ and $V^{\prime}$.
Proposition 5.4.
Let $\mathcal{E_{V}}$ denote a relation on the set of all controlled stochastic automata such that $(V_{1},V_{2})\in\mathcal{E_{V}}$ if and only if $V_{1}$ and $V_{2}$ are equivalent controlled stochastic automata in the sense of Definition  5.3. Then $\mathcal{E_{V}}$ will be an equivalence relation.
Operational Semantics
Definition 5.5.
Let $N=(P,IA,TA,CA,IG,OG,IR,IOR,TOR,IP,\overline{F},\overline{\rho},\overline{\Pi})$ be a controlled stochastic activity network with an initial marking $\mu_{0}$. $(N,\mu_{0})$ is said to realize a controlled stochastic automaton $V=(Q,A,C,P,Q_{0,},F,\rho,\Pi)$ where:
  • $U=(Q,A,C,P,Q_{0})$ is a controlled probabilistic automaton realized by $(L,\mu_{0})$, where $L=(P,IA,TA,$ $CA,IG,OG,IR,IOR,TOR,IP)$,
  • for any $q\in Q$ and $a\in A$, $F(.|q,a)=\overline{F}(.|q,a),$ $\rho(q,a)=\overline{\rho}(q,a)$ and $\Pi(q,a)=\overline{\Pi}(q,a)$.
A notion of equivalence for controlled stochastic activity networks may now be given as follows.
Definition 5.6.
Two controlled stochastic activity networks are equivalent if they realize equivalent controlled stochastic automata.
Proposition 5.7.
Let $\mathcal{E_{N}}$ denote a relation on the set of all controlled stochastic activity networks such that $(N_{1},N_{2})\in\mathcal{E_{N}}$ if and only if $N_{1}$ and $N_{2}$ are equivalent controlled stochastic activity networks in the sense of Definition  5.6. Then $\mathcal{E_{N}}$ will be an equivalence relation.
Policy Types
Given a controlled stochastic automaton $V=(Q,A,C,P,Q_{0,},F,\rho,\Pi)$. The model can be viewed to evolve over continuous time, where the completion of an activity in a state triggers a control decision that determines the next state. Let $X(t)$, $Y(t)$, and $Z(t)$ represent the state, control action, and activity completed (if any) at time $t$, respectively. ($Z(t)=\emptyset$ if no activity completes at time $t$.) Let $\{X(t^{\prime});0\leq t^{\prime}<t,t^{\prime}\in\mathcal{R}_{+}\}. H(t)H(t)$ is called the history of $V$ before $t$. A (history-depenent) policy for $V$ is a set $\pi=\{\pi(t);t\in\mathcal{R}_{+}\}$ where, for any time $t,$ $\pi(t)(.|X(t),H(t))$ is a conditional probability measure over $C$. Suppose activity a completes immediately before time $t$. Then, at time $t$, policy $\pi$ chooses a control action $c$ with probability $\pi(t)(c|X(t),H(t))$. $\pi$ is said to be a Markov policy if $\pi(t)(.|X(t),H(t))$ is independent of $H(t)$ for any time $t$. A Markov policy $\pi$ is said to be a memoryless policy if $\pi(t)(.|X(t))$ is further independent of $t.$. It immediately follows that any memoryless policy $\pi$ may be represented as a set $\pi=\{\pi(.|q);q\in Q\}$where $\pi(.|q)$ is a conditional probability over $C$ for any $q\in Q$
Having introduced the behavioral and structural semantics of CSAs, we now turn to the quantitative aspects of decision-making. In order to analyze performance and synthesize optimal control strategies, we must augment CSAs with a reward framework that assigns costs or benefits to states, control actions, and activity completions. This reward structure serves as the foundation for comparing policies, establishing optimality criteria, and studying complexity results across both discrete- and continuous-time settings.
Reward Model and PAC-Optimality
We impose a reward structure $r=(r^{\prime},r^{\prime\prime})$, where $r^{\prime}$ is a bounded Borel measurable function
$r^{\prime}:Q\times C\to\mathbb{R}_{+},$
called the reward rate function, and $r^{\prime\prime}$ is a bounded Borel measurable function
$r^{\prime\prime}:Q\times A\times C\times Q\to\mathbb{R}_{+},$
called the activity reward function. Let $D_{t}$ denote the number of activity completions by time $t$ and $t^{+}$ a time immediately after $t$. The reward incurred during a period $\tau^{\prime},\tau]$, for $\tau^{\prime},\tau\in\mathbb{R}_{+}$ with $\tau^{\prime}\leq\tau$, is defined by the random variable
$R=\int_{\tau^{\prime}}^{\tau}r^{\prime}(X(t),Y(t))\,dt+\int_{\tau^{\prime}}^{\tau}r^{\prime\prime}(X(t),Z(t),Y(t),X(t^{+}))\,dD_{t}.$ (5.1)
The variable $R$ is well-defined under the boundedness and measurability conditions above.
To compare rewards, we use the stochastic ordering $\leq_{st}$ on random variables: for $X,Y$, we write $X\leq_{st}Y$ iff $\Pr[X>x]\leq\Pr[Y>x]$ for all $x\in\mathbb{R}_{+}$.
Let $V=(Q,A,C,P,Q_{0},F,\rho,\Pi)$ be a controlled stochastic automaton (CSA) with reward structure $r$, and let $\mathcal{P}$ denote a class of admissible policies. For $\tau^{\prime},\tau\in\mathbb{R}_{+}$ with $\tau^{\prime}\leq\tau$, $\pi\in\mathcal{P}$, and $q\in Q$, define
$R_{\tau^{\prime}}^{\tau}(\pi\mid q)\equiv\text{the reward accrued during }[\tau^{\prime},\tau]\text{ given }X(\tau^{\prime})=q.$
A policy $\pi^{*}\in\mathcal{P}$ is called an optimal policy for $(V,r)$ over $\tau^{\prime},\tau]$ if for all $\pi\in\mathcal{P}$ and $q\in Q$,
$R_{\tau^{\prime}}^{\tau}(\pi\mid q)\leq_{st}R_{\tau^{\prime}}^{\tau}(\pi^{*}\mid q).$
Such an optimal policy, however, is a strong requirement and may not always exist.
In many realistic applications, the exact computation of optimal policies for continuous-time decision processes are infeasible due to the size and complexity of the underlying state space or stochastic dynamics. This motivates the use of Probably Approximately Correct (PAC) methods [valiant1984theory], which provide theoretical guarantees of $\varepsilon$-optimality with high probability using sampling- or learning-based procedures. Such guarantees make PAC algorithms particularly attractive for systems in which exact analysis is intractable, but near-optimal performance can be learned efficiently.
Remark 5.8.
For any $\varepsilon,\delta>0$, there exist PAC reinforcement learning algorithms that return a policy $\pi$ such that, with probability at least $1-\delta$, the expected performance of $\pi$ is within $\varepsilon$ of the optimal value. Foundational results for finite-state MDPs were established in [kearns2002near, strehl2009reinforcement], with later refinements providing tight bounds for discounted MDPs [lattimore2013pac]. These methods can be adapted to continuous-time decision processes through uniformization or related discretization techniques, extending PAC guarantees to stochastic systems modeled by Controlled SANs.
Integrating PAC guarantees into the framework of Controlled SANs provides a principled way to synthesize near-optimal policies when traditional analytic optimization methods are infeasible. Unlike exact approaches, PAC algorithms offer performance bounds that scales with problem complexity, enabling learning-based control for high-dimensional or partially specified systems [SuttonBarto1998]. This connects the classical theory of optimal and $\varepsilon$-optimal policies [Howard71, puterman1994markov] with modern sample-efficient learning frameworks [valiant1984theory, kearns2002near, strehl2009reinforcement, lattimore2013pac]. This bridge is particularly relevant for AI-driven applications such as autonomous systems, adaptive scheduling, and dependable real-time control, where rigorous performance guarantees and learning-based adaptability must coexist within a unified formal framework.
The notion of stochastic equivalence in Definition 5.3 allows us to compare the behavior of different random processes realized by controlled stochastic automata. In many practical applications, however, it is crucial to identify when such an automaton exhibits the simpler Markovian behavior, as this enables more tractable analysis and controller synthesis. We next characterize conditions under which a controlled stochastic automaton can be classified as Markovian. To compare formally the behavior of various models, we need the following concepts.
Definition 5.9.
The state process of a model is a random process $\{X(t);t\in{\mathcal{R}_{+}}\}$ where $X(t)$ denotes the state of the model at time $t$.
Definition 5.10.
Let $X=\{X(t);t\in{\mathcal{R}_{+}}\}$ and $X^{\prime}=\{X^{\prime}(t);t\in{\mathcal{R}_{+}}\}$ be two random processes with the set of states $Q$ and $Q^{\prime}$, respectively. $X$ and $X^{\prime}$ are said to be stochastically equivalent if there exists a symmetric binary relation $\gamma$ on $Q\cup Q^{\prime}$ such that:
  • for any $q\in Q$, there exists $q^{\prime}\in Q^{\prime}$ sich that, $(q,q^{\prime})\in\gamma$; also for any $q^{\prime}\in Q^{\prime}$, there exists $q\in Q$ such that $(q^{\prime},q)\in\gamma$,
  • for any $t_{i}\in[0,\;\infty)$, $Q_{i}\subseteq Q$, and $Q^{\prime}_{i}\subseteq Q^{\prime}$, such that for any $q\in Q_{i}$, there exists $q^{\prime}\in Q^{\prime}_{i}$ such that $(q,q^{\prime})\in\gamma$; also and for any $q^{\prime}\in Q^{\prime}_{i}$, there exists $q\in Q_{i}$ such that $(q^{\prime},q)\in\gamma$, $i=0,\ldots,n$, $n\in{\mathcal{N}}$,
    $p[X(t_{i})\in Q_{i};i=0,\ldots,n]=p[X^{\prime}(t_{i})\in Q^{\prime}_{i};i=0,\ldots,n].$
$X$ and $X^{\prime}$ are stochastically isomorphic (equal) if $\gamma$ is a bijection (an equality).
The ability to synthesize policies depends heavily on the underlying process structure. Certain CSAs exhibit Markovian dynamics, which allow for more tractable analysis and standard solution techniques. Identifying and exploiting such structure bridges our framework with classical models such as Continuous-Time Markov Decision Processes (CTMDPs), enabling the application of established algorithms while retaining the expressive power of CSAs.
Controlled Markovian Automata
Proposition 5.11.
Let $V=(Q,A,C,P,Q_{0},F,\rho,\Pi)$ be a controlled stochastic automaton. The state process of $V$ is a Markov process if for any policy $\pi$ and activity $a$ which is enabled in a state $q$, and any state $q_{ac}$ in which $a$ was last activated before being enabled in $q$, and $\tau\geq 0$,
$F(\tau|q_{ac},a)=1-e^{-\alpha(q,a)\tau}$
where $\alpha(q,a)$ is a positive real number, referred to as the activity time rate of a in q. Then, $V$ is called to be Markovian.
This characterization enables us to represent Markovian controlled stochastic automata more succinctly. In particular, we can exploit their structure to define a compact formalism, which will facilitate the subsequent development of value iteration and policy synthesis. We formalize this representation in the following definition, where we present a compact representation of a Markovian controlled stochastic automaton.
Definition 5.12.
A controlled Markovian automaton (CMA) is a 6-tuple $W=(Q,A,C,P,Q_{0},\sigma)$ where:
  • $(Q,A,C,P,Q_{0})$ is a controlled probabilistic automaton,
  • $\sigma:Q\times A\to\mathcal{R}_{+}$ is the actvity rate function,
Proposition 5.13.
Let $V=(Q,A,C,P,Q_{0},F,\rho,\Pi)$ be a Markovian controlled stochastic automaton with a policy $\pi_{V}$. There exists a controlled Markovian automaton $W=(Q,A,C,P,Q_{0},\sigma)$, where
$\sigma(q,a)=\rho(q,a)\alpha(q,a)$
with a policy $\pi_{W}$ such that the state process of $V$ and $W$ are stochastically isomorphic under policies $\pi_{V}$ and $\pi_{W}$, respectively. $V$ is said to be represented by $W$.
Definition 5.14.
A controlled Markovian activity network is a controlled stochastic activity network that, in any initial marking, realizes a Markovian controlled stochastic automaton. Using Proposition  5.13 above, a controlled Markovian network is, equivalently, represented by a controlled Markovian automaton.
Continuous-Time Markov Decision Process (CTMDP)
We now introduce a popular and general framework commonly used in most theoretical decision-making problems that evolve over continuous time, and its relation to our proposed models is as follows.
Definition 5.15.
A continuous-time Markov decision process (CTMDP) is defined as a 5-tuple $D=(Q,C,Q_{0},\lambda,P)$, where:
  • Q is a countable set of states,
  • C is a finite set of control actions,
  • $Q_{0}$ is the initial state distribution where $\sum_{q\in Q}Q_{0}(q)=1$,
  • $\lambda:Q\times C\to\mathcal{R}_{+}$ is the transition rate function,
  • $P:Q\times C\times Q\to[0,1]$ is the transition probability function, where for any $q\in Q$ and $c\in C$,
    $\sum_{q^{\prime}\in Q}P(q,c,q^{\prime})=1.$
Definition 5.16.
Let $W=(Q,A,C,P,Q_{0},\sigma)$ be a controlled Markovian automaton (CMA). Then, $W$ is said to be represented by a continuous-time Markov decision process (CTMDP) $D=(Q,C,Q_{0},\lambda,P^{\prime})$ defined as:
$\lambda(q,c)=\sum_{a\in A}\sigma(q,a)\sum_{q^{\prime}\in Q}P(q,a,c,q^{\prime})$
$P^{\prime}(q,c,q^{\prime})=\frac{\sum_{a\in A}\sigma(q,a)P(q,a,c,q^{\prime})}{\sum_{a\in A}\sigma(q,a)\sum_{q^{\prime}\in Q}P(q,a,c,q^{\prime})}$
In the above model representation, the activity component becomes implicit, and the CTMDP models the same transition dynamics by summarizing the rate and probability information from the CMA. Thus, a CTMDP model can be viewed as an abstract representation of a CMA model. More formally, we have:
Proposition 5.17.
Let $W=(Q,A,C,P,Q_{0},\sigma)$ and $D=(Q,C,Q_{0},\lambda,P^{\prime})$ be a controlled Markovian automaton (CMA) and a continuous-time Markov decision process (CTMDP), respectively, as defined in Definition  5.16 above. Then, there exist policies $\pi_{W}$ and $\pi_{D}$ in $W$ and $D$, respectively, such that the state processes of $W$ and $D$ under policies $\pi_{W}$ and $\pi_{D}$, respectively, will be isomorphic.
Having established the correspondence between Controlled SANs and continuous-time Markov decision processes (CTMDPs) via controlled Markovian automata, we now shift our focus to the computational aspects of controller synthesis. While the preceding definitions clarified how CSA-derived systems can be abstracted as CTMDPs, practical applications require understanding when optimal policies exist and how they can be computed efficiently. We therefore analyze the complexity of finding optimal policies under standard quantitative objectives, beginning with the case of finite-state CTMDPs, where classical algorithmic techniques can be leveraged directly.
Optimal Policies for Finite-State CTMDPs
We examine the synthesis of optimal policies for continuous-time Markov decision processes (CTMDPs) that arise from Controlled SANs when all activities are exponentially timed, assuming finite state and control action sets. By the bounded reward assumptions introduced in the Reward Model and PAC-Optimality Subsection earlier, and standard results for finite-state CTMDPs (see, e.g., [puterman1994markov, feinberg2002handbook]), an optimal stationary deterministic policy is guaranteed to exist. The computational complexity of finding such a policy depends on the specific quantitative objective, as illustrated in the following property:
Proposition 5.18.
Consider a finite-state continuous-time Markov decision process (CTMDP) induced by a Controlled SAN in which all activities are exponentially distributed, and assume bounded reward functions. Then:
  1. 1.
    For expected total reward, there always exists an optimal memoryless deterministic policy, and the corresponding value function can be computed in polynomial time.
  2. 2.
    For discounted reward, an optimal memoryless deterministic policy exists, and the value function can be obtained by solving a linear system in polynomial time.
  3. 3.
    For time-bounded reachability or reward, the problem is PSPACE-complete in general. Optimal policies may require history dependence, and existing numerical algorithms exhibit exponential dependence on the required precision.
Proof Sketch.
The existence of optimal stationary deterministic policies for expected total and discounted rewards in finite-state CTMDPs follows from standard MDP optimality results. For expected total reward, the optimization problem can be formulated as a linear program, guaranteeing a memoryless deterministic optimal policy and polynomial-time solvability. For a discounted reward, the Bellman equations yield a linear system that can be efficiently solved using standard numerical linear algebra methods [puterman1994markov].
The time-bounded reachability (or reward) objective is substantially harder: computing optimal values involves solving systems of ODEs or using uniformization/discretization techniques. The associated decision problems are PSPACE-complete, and optimal strategies may require history-dependent policies [courcoubetis1995complexity]. ∎
Remark 5.19.
The tractability gap between the discounted/total reward and the time-bounded objectives reflects the fundamental increase in computational complexity introduced by finite-horizon constraints in continuous time. While polynomial-time methods suffice for the first two cases, practical algorithms for the third objective rely on numerical approximations with potentially high computational cost.
Table 3 summarizes the computational complexity and policy characteristics for optimal control in finite-state CTMDPs under the three standard objectives discussed in Proposition 5.18. These results establish that when Controlled SANs yield CTMDPs through exponentially timed activities, existing algorithmic techniques from the CTMDP literature can be directly applied for controller synthesis in safety-critical applications. Such reductions form the basis for extending these methods to more general CSA models, where non-exponential activities lead beyond the classical CTMDP framework.
Table 3: Complexity and policy characteristics for optimal control in finite-state CTMDPs, categorized by quantitative objective.
Objective Optimal Policy Type Complexity Class Solution Method
Expected Total Reward Memoryless deterministic Polynomial time Linear programming [puterman1994markov]
Discounted Reward Memoryless deterministic Polynomial time Linear system or value iteration [puterman1994markov]
Time-Bounded Reachability / Reward May require history dependence PSPACE-complete ODE solving or uniformization [courcoubetis1995complexity]
In summary, for finite-state CTMDPs, both expected total and discounted reward objectives admit efficient algorithms and memoryless optimal policies, whereas time-bounded objectives lead to significantly higher complexity and may require history-dependent strategies. These results provide a clear computational boundary between classical tractable control objectives and those that require more sophisticated approximation or synthesis techniques.
The finite-state results provide the foundation for handling Controlled SANs with unbounded or continuous state spaces, where CTMDPs can no longer be solved directly. We now turn to the synthesis of optimal policies for infinite-state CTMDPs, highlighting the challenges and methods for approximating $\epsilon$-optimal controllers in this broader setting.
The results for finite-state models, summarized in the preceding subsection, provide a foundation for analyzing more expressive Controlled SAN instances where the underlying CTMDP may exhibit an infinite, though countable, state space. Such cases arise naturally from complex systems with unbounded counters, queues, or dynamically generated processes. In these settings, standard CTMDP algorithms no longer apply directly, and both the existence and approximation of optimal policies become more challenging. We now extend the discussion to infinite-state CTMDPs, identifying the conditions for policy existence and techniques for constructing $\epsilon$-optimal controllers.
Optimal Policies for Infinite-State CTMDPs
In many applications, the state space of a Controlled SAN—when interpreted as a continuous-time Markov decision process (CTMDP)—may be infinite but countable. We now address the question of whether optimal policies exist in this more general setting.
The results for finite-state models, summarized in the previous subsection, provide a foundation for analyzing more expressive Controlled SAN instances where the underlying CTMDP may exhibit an infinite, though countable, state space. Such cases arise naturally from complex systems with unbounded counters, queues, or dynamically generated processes. In these settings, standard CTMDP algorithms no longer apply directly, and both the existence and approximation of optimal policies become more challenging. We now extend the discussion to infinite-state CTMDPs, identifying the conditions for policy existence and techniques for constructing $\epsilon$-optimal controllers.
Proposition 5.20.
Let $D=(Q,C,Q_{0},\lambda,P)$ be a CTMDP derived from a Controlled SAN, where $Q$ is a countably infinite set. Recall the reward structure imposed in our reward model mentioned earlier in this section, namely, $r=(r^{\prime},r^{\prime\prime})$, where $r^{\prime}$, the reward rate function, and $r^{\prime\prime}$, the activity reward function, are both assumed to be bounded Borel measurable functions. Our goal is to maximize the expected total accumulated reward. Then, we have:
  1. 1.
    The value function is well-defined and bounded;
  2. 2.
    There exists an optimal policy attaining the supremum reward;
  3. 3.
    If $r^{\prime}$ is continuous (in a measurable structure) and $r^{\prime\prime}$ is weakly continuous, then we can have a stationary deterministic policy.
If the reward is discounted by a factor $\beta>0$, the same conclusions hold under even weaker assumptions, and the optimal policy can be computed via value iteration or fixed-point techniques.
Proof Sketch.
The proof follows standard arguments for infinite-state MDPs extended to the continuous-time setting by uniformization, combined with bounded reward assumptions and Banach fixed-point arguments. See [puterman1994markov, HernandezLerma1996] for classical existence results for optimal stationary policies. ∎
Remark 5.21.
If the reward function $r$ is unbounded or if the system admits Zeno behavior (i.e., an infinite number of transitions in finite time), then the expected reward may diverge, and an optimal policy may not exist. Moreover, for time-bounded objectives, only $\varepsilon$-optimal policies are generally guaranteed under standard measurability and compactness assumptions.
Having established the existence and computability of optimal policies for both finite- and infinite-state CTMDPs derived from Controlled SANs, we now examine the expressive power of our modeling framework itself. This subsection synthesizes prior results on policy structures, reward semantics, and complexity to show how Controlled Stochastic Automata (CSAs) relate to, subsume, or differ from classical models such as CTMDPs, GSMPs, and other stochastic systems. By clarifying these relationships, we provide a comprehensive perspective on where CSA fits in the hierarchy of stochastic models and what this implies for policy synthesis.
Expressiveness and Optimality of CSA and CTMDP Models
This subsection consolidates the main results on Controlled Stochastic Automata (CSA), connecting the framework introduced earlier in Subsection Policy Types and Subsection Reward Model and PAC-Optimality to the analysis of expressiveness and optimality. We first establish the relationship of CSA to classical models such as CTMDPs and GSMPs, clarifying its expressiveness boundaries. Building on the policy structures and reward accumulation mechanisms defined in Subsections Policy Types and Reward Model and PAC-Optimality, we then present a unified treatment of policy synthesis and value iteration, covering both the general case with arbitrary timing distributions (via GSMP/SMDP embedding) and the special case where exponential timing allows a direct CTMDP reduction.
Proposition 5.22.
Controlled Stochastic Automata (CSA) with general timing distributions strictly generalize classical continuous-time Markov decision processes (CTMDPs) and generalized semi-Markov processes (GSMPs). Specifically:
  1. 1.
    Any GSMP can be represented as a CSA by encoding each enabled activity’s clock and transition kernel within the CSA’s timing and control structure.
  2. 2.
    CSA subsumes CTMDPs whenever all timing distributions are exponential, since the memoryless property removes the need to track residual holding times.
  3. 3.
    CSA cannot, in general, embed stochastic hybrid systems (SHS) or piecewise deterministic Markov processes (PDMPs), because their continuous state dynamics and differential flows cannot be expressed within CSA’s discrete-event semantics.
Sketch of Proof.
For (1), a CSA with general (possibly non-exponential) sojourn times can be mapped to a GSMP by associating each enabled activity $a$ under control $c$ with its timing distribution and transition kernel. The GSMP state, augmented with residual timers, captures exactly the CSA’s event-driven dynamics [Matthes1962, Schassberger1978].
For (2), when all sojourn times are exponential, residual-time information becomes unnecessary due to the memoryless property, and the CSA reduces to a CTMDP. Standard uniformization then yields a discrete-time MDP semantics, enabling classical CTMDP solution techniques.
For (3), SHS and PDMPs include continuous state spaces with deterministic flows governed by differential equations. CSA, being a purely discrete-event formalism with stochastic timing, cannot capture such continuous flows, though it may approximate event-driven subclasses of these models [Cassandras2008, Tomlin1998, Davis1984]. ∎
Proposition 5.23.
Let $V$ be a Controlled Stochastic Automaton (CSA) with bounded reward rates, non-explosive dynamics, and general (not necessarily exponential) timing distributions, as specified under the policy and reward semantics in Subsections Policy Types, and Reward Model and PAC-Optimality. For any discount factor $\beta>0$:
  1. 1.
    There exists a sequence of discretized semi-Markov decision processes (SMDPs) approximating $V$, whose optimal value functions converge uniformly to the supremum expected discounted reward achievable under all admissible CSA policies.
  2. 2.
    The Bellman operator $T$ associated with each discretized model is a contraction (with modulus $e^{-\beta\Delta}$ for discretization step $\Delta$), so value iteration $V_{k+1}:=TV_{k}$ converges geometrically to the optimal value function $V^{*}_{\Delta}$.
  3. 3.
    For any $\varepsilon>0$, a sufficiently fine discretization yields a stationary (possibly randomized) policy that is $\varepsilon$-optimal for $V$.
  4. 4.
    When all timing distributions are exponential, the CSA reduces to a CTMDP, which can be uniformized into a discrete-time MDP. In this case, classical CTMDP value iteration applies directly and converges to the optimal value function without state augmentation.
Sketch of Proof. For general timing, we embed $V$ into a GSMP/SMDP by extending each state with residual times of all enabled timed activities, obtaining a Markovian extended state space. Discretizing time with a step size $\Delta>0$, we form the embedded chain, which captures event completions and residual-time evolution. This yields a discrete-time decision process with bounded one-step rewards.
The Bellman operator $T$ for the discretized model is a contraction with modulus $\gamma=e^{-\beta\Delta}$, so the sequence $V_{k+1}:=TV_{k}$ converges geometrically to the unique fixed point $V^{*}_{\Delta}$. Standard error bounds for discretization are well established, both in the SMDP setting (cf. Howard [Howard71, Chapter 15]) and in the MDP setting (cf. Puterman [puterman1994markov, Chapter 7]). These results imply that $V^{*}_{\Delta}$ converges uniformly to the supremum discounted reward $V^{*}$ for the CSA as $\Delta\to 0$. Thus, for any $\varepsilon>0$, a sufficiently fine discretization yields a stationary policy that is $\varepsilon$-optimal for $V$.
Direct uniformization is not applicable unless all sojourn times are exponential. If they are, residual timers are unnecessary by the memoryless property, and $V$ is equivalent to a CTMDP. Standard uniformization transforms this CTMDP into a discrete-time MDP, allowing for the application of classical value iteration, which converges by contraction arguments. ∎
Remark 5.24.
While the above results guarantee the existence of $\varepsilon$-optimal policies and convergence of value iteration, the computational complexity depends heavily on the timing structure. For CTMDPs (exponential case), uniformization leads to tractable dynamic programming with existing solvers. For general CSAs, residual-time augmentation can cause significant state-space growth, and discretization introduces additional overhead. Practical implementations typically rely on truncation, state aggregation, or approximation heuristics to manage this complexity while preserving near-optimality guarantees.
In general, a Controlled Stochastic Automaton (CSA) may include timed activities with arbitrary, non-exponential distributions. Unlike the special case of exponential timing, where the process reduces to a continuous-time Markov decision process (CTMDP) and can be directly uniformized into a discrete-time MDP, non-exponential sojourn times break the memoryless property and preclude such a reduction. To handle the full CSA model, we instead embed it into a generalized semi-Markov process (GSMP) or semi-Markov decision process (SMDP) by augmenting the state with residual timers. This construction preserves the underlying dynamics in a Markovian extended state space and enables discretization-based dynamic programming.
These results collectively show that, while CSA and CTMDP models can exhibit structural complexities that preclude exact optimal policies, $\epsilon$-optimal policies always exist and can be systematically computed. Combined with the Reward Model framework, they ensure that the accumulated reward functional $R_{\tau^{\prime}}^{\tau}(\pi\mid q)$ can be approximated within any tolerance $\epsilon$, allowing robust analysis and synthesis even in complex settings.
6 Conclusion
This paper has introduced Controlled Stochastic Activity Networks (Controlled SANs) as a formal, automata-theoretic framework that unifies control, nondeterminism, probability, and stochastic timing within a single modeling language. By developing a layered semantic hierarchy—from controlled automata to probabilistic and stochastic extensions—we have established a rigorous foundation for representing and analyzing distributed real-time systems operating under uncertainty.
Our results formally connect Controlled SANs with classical models such as CTMDPs, GSMPs, and controlled probabilistic automata, while also introducing a structured taxonomy of control policies and corresponding language hierarchies. These theoretical developments enable precise reasoning about expressiveness, equivalence, and compositionality.
A distinctive aspect of this work is its compatibility with AI-driven control synthesis. The framework naturally supports PAC-based reasoning, reinforcement-learning approaches, and other algorithmic techniques for scalable decision making under uncertainty. This bridges classical formal verification with modern AI methods, creating new opportunities for practical verification and adaptive control in complex systems.
Future work will focus on integrating temporal-logic model checking, controller synthesis, and PAC-optimization methods into automated tools, as well as exploring applications in safety-critical domains such as autonomous systems, dependable AI, and cyber-physical infrastructures. By combining rigorous formal semantics with learning-based control, Controlled SANs provide a mathematically principled foundation for the next generation of verifiable intelligent systems.
AI Assistance Statement
During the preparation of this work, the author used ChatGPT (OpenAI) to assist with language polishing, technical phrasing, and improving the clarity of exposition. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
Appendix A Proofs
Proof of Theorem  3.13
Proof.
Consider the class of activity networks [movaghar2001stochastic] that have only instantaneous activities, standard gates, and a special type of gates called “inhibitor” gates. An inhibitor gate is an input gate with an enabling predicate $g$ and an identity function such that $g(x)=true$ iff $x=0$. The above class of activity networks corresponds to the class of extended Petri nets [Peterson1981]. Using the notion of the register machines of Shepardson and Sturgis [Shepardson1963] or the program machines of Minsky [Minsky1967], it can easily be shown that an extended Petri net can simulate a Turing machine [Hack1975, Peterson1981]. Following a similar approach, one can show that an extended Petri net can, indeed, simulate a nondeterministic Turing machine [Sipser2013] and, hence, can represent any computable relation.
It follows that an activity network with some instantaneous activities, standard gates, and inhibitor gates is, likewise, able to simulate a nondeterministic Turing machine and, hence, can represent any computable relation. This latter result will be used below to prove the theorem.
Let $S=(Q,A,C,\rightarrow,Q_{0})$ be a computable controlled automaton. Without loss of generality, let $Q={\mathcal{N}}$. Define:
  • for $a\in A$ and $c\in C$, $R_{ac}=\{(i,j);i,j\in{\mathcal{N}},i\stackrel{{\scriptstyle a,c}}{{\rightarrow}}j\}$
  • for $a\in A$, $G_{a}:{\mathcal{N}}\longrightarrow\{true,false\}$, where $G_{a}(i)=true$ iff there exists $c\in C$ and $j\in{\mathcal{N}}$ such that $i\stackrel{{\scriptstyle a,c}}{{\rightarrow}}j$,
  • $R_{Q_{0}}=\{(1,i);i\in Q_{0}\}$.
Since $S$ is a computable controlled automaton, $G_{a},R_{ac}$, for $a\in A$ and $c\in C$, and $R_{Q_{0}}$ above will also be computable relations, which means that some nondeterministic Turing machines can represent them. Accordingly, using the result mentioned earlier, $G_{a},R_{ac}$, for $a\in A$ and $c\in C$, and $R_{Q_{0}}$ can be simulated by some activity networks $K_{G_{a}},K_{R_{ac}}$, for $a\in A$ and $c\in C$, and $K_{Q_{0}}$, respectively, which have only some instantaneous activities, standard gates and inhibitor gates.
Now, consider a controlled activity network $K$ with the set of timed activities $A$ and the set of contol actions $C$ such that, for any $a\in A$ and $c\in C$, $K$ includes an controlled activity subnetwork as depicted in Figure 5. In Figure 2, $K_{G_{a}}$ and $K_{R_{ac}}$ represent some controlled activity networks which simulate $G_{a}$ and $R_{ac}$, respectively. $P_{1a}$, $P_{2a}$, $P_{3ac}$, and $P_{M}$ are places such that when $P_{1a}$ is empty all instantaneous activities of $K_{G_{a}}$ are disabled and when $P_{3ac}$ is empty all instatntaneous activities of $K_{R_{ac}}$ are disabled. Initially, $P_{1a}$ has a token but $P_{2a}$ and $P_{3ac}$ are empty. As soon as $P_{S}$ acquires a marking $x$ such that $G_{a}(x)=true$, $K_{G_{a}}$ starts execution and after a finite number of instantaneous activity completions $P_{1a}$ loses a token, $P_{2a}$ gains one, and the marking of $P_{S}$ remains the same. When $P_{3ac}$ obtains a token, $K_{R_{ac}}$ also begins execution and after a finite number of instantaneous activity completions $P_{3a}$ loses a token, $P_{1a}$ gains one, and the marking of $P_{S}$ changes from $x$ to $y$ where $(x,y)\in R_{ac}$. $K$ also includes a controlled activity subnetwork as depicted in Figure 6. In Figure 3, $K_{Q_{0}}$ represents a controlled activity network which simulates $R_{Q_{0}}$. $P_{0}$ is a place with an initial marking of 1. Initially, $P_{S}$ is empty and $K_{Q_{0}}$ starts execution. After a finite number of instantaneous activity completions $P_{0}$ loses a token, the marking of $P_{S}$ is set to $x$ where $(1,x)\in R_{Q_{0}}$, and all activities of $K_{Q_{0}}$ are disabled. Suppose $K$ includes no controlled activity subnetwork other than those depicted in Figure 5 and Figure 6. It can be shown that the controlled activity network $K$ defined above, indeed, realizes a controlled automaton which is isomorphic to $S$ mentioned earlier, and the proof is complete. ∎
Figure 1: An example of a simple controlled activity network with a marking.
Figure 2: Marking of the model after $T1$ completes via control action $c1$ in the model of Figure 1.
Figure 3: Marking of the model after $I2$ completes in the model of Figure 2.
Figure 4: Marking of the model after $I1$ completes in the model of Figure 3.
$K_{G_{a}}$$P_{2a}$  $acP_{3ac}$$K_{R_{ac}}$$P_{1a}$$P_{S}$Figure 5. A controlled activity subnetwork of KKcorresponding to an activity $a$ and control action $c$ of $M$.
$P_{0}$$K_{Q_{0}}$$P_{S}$Figure 6. A controlled activity subnetworkof $K$ corresponding to $Q_{0}$.